### v1

# Equilibrium Portfolio Selection under Utility-Variance Analysis of Log Returns in Incomplete Markets

Abstract
This paper investigates a time-inconsistent portfolio selection problem in the incomplete market model, integrating expected utility maximization with risk control. The objective functional balances the expected utility and variance on log returns, giving rise to time inconsistency and motivating the search of a time-consistent equilibrium strategy. We characterize the equilibrium via a coupled quadratic backward stochastic differential equation (BSDE) system and establish the existence theory in two special cases: (i) two Brownian motions driven the price dynamics and the factor process are independent with $\rho=0$; (ii) the trading strategy is constrained to be bounded. For the general case with correlation coefficient $\rho\neq 0$, we introduce the notion of an approximate time-consistent equilibrium. By employing the solution structure from the equilibrium in the case $\rho=0$, we can construct an approximate time-consistent equilibrium in the general case with an error of order $O(\rho^{2})$. Numerical examples and financial insights based on deep learning algorithms are also presented.
Keywords: Time inconsistent control, time-consistent equilibrium, quadratic BSDE system, approximate time-consistent equilibrium
1 Introduction
The optimal portfolio management problem has always been a core research topic in quantitative finance, traditionally approached via the mean-variance (MV) criterion proposed by Markowitz_1952 and the expected utility (EU) theory developed in merton_1969; MERTON1971373. The classical mean–variance criterion captures the trade-off between the expected return and the risk (variance). In contrary, the utility maximization framework incorporates the investor’s risk aversion level into decision making without concerning the risk level of terminal wealth.
One natural extension is to integrate the risk management into the utility maximization problem, dictating the optimal trading strategy at a reasonable risk level. The research in this direction has gained an upsurge of attention recently. As a pioneer attempt, Bsask_2001 study the utility maximization on portfolio and consumption plans by imposing the Value-at-Risk (VaR) constraints on the terminal wealth to encode the agent’s concern on the risk level. Wong_2017 examine a utility–deviation–risk portfolio selection problem by reformulating it as an equivalent nonlinear moment system, where the objective combines expected utility with a deviation-based risk measure, and derives the optimal strategy in a Black-Scholes model. Bensouusan_2022 recently combine the utility maximization problem with the variance minimization of terminal wealth, and recasts the problem as a mean field-type control (MFC) problem using a coupled system of Hamilton–Jacobi–Bellman and Fokker–Planck equation, and characterizes the optimal consumption and trading strategies in the mean-field context.
Our paper adopts an integrated formulation that aims to optimize the trade-off between the expected utility and the variance of portfolio in incomplete market models. Unlike the setting that focuses on the capital amount (see, e.g., Wong_2017 and Bensouusan_2022), we target at the log returns of the portfolio. This choice follows a growing line of research on portfolio selection (see dai_dynamic_2021, Peng_2023, Guan2025, and references therein). Specifically, the objective functional in our paper at time $t$ is defined as
$J(t,\pi)=\mathbb{E}_{t}\left[U\left(\log\left(\frac{W^{\pi}_{T}}{W^{\pi}_{t}}\right)\right)\right]-\frac{\gamma_{t}}{2}\textup{Var}_{t}\left[\log\left(\frac{W^{\pi}_{T}}{W^{\pi}_{t}}\right)\right].$ (1.1)
Nevertheless, the log returns and the variance term in the objective functional render the problem time-inconsistent. That is, a strategy being optimal today may no longer remain optimal at future dates. Strotz1955 classifies agents facing time-inconsistency into three types. The naive agent always re-optimizes the current objective, ignoring time-inconsistency and thus constantly changing strategies. The precommitted agent fixes an initial optimal plan, knowing it will later become suboptimal; the plan remains time-inconsistent since it is only optimal at the initial time. A sophisticated agent cannot precommit but anticipates future deviations, viewing her future selves as strategic counterparts and seeking a time-consistent equilibrium strategy that each self optimally follows—namely, the intrapersonal equilibrium.
Contrary to Wong_2017 and Bensouusan_2022, who focus on the precommitted strategy, the present paper aims to characterize a time-consistent equilibrium strategy by a sophisticated agent under game theoretical thinking with future selves, whose precise definition in continuous time was first proposed in Ekeland2006noncommitmentsubgameperfectequilibrium [^1] Since then, a strand of literature on time-inconsistent control problems in continuous-time setting has emerged, see e.g. Basak_dynamic_2010, hu_time-inconsistent_2012, hu_time-inconsistent_2017, Bjork_2014, dai_dynamic_2021 for MV problems, Ekeland2006noncommitmentsubgameperfectequilibrium, Ekeland_2007, Hamaguchi_2021 for non-exponential discounting problems.. Moreover, we consider the incomplete market model, where the financial market consists of a risk-free asset and a single risky asset [^2] For simplicity, we focus on one risky asset, as the analysis can be extended directly to multiple risky assets.. The wealth process is driven by a standard Brownian motion $B$, while the asset prices are influenced by an exogenous factor process driven by another Brownian motion $\bar{B}$. The correlation coefficient of these two Brownian motions is denoted by $\rho$. Similar to Hamaguchi_2021 and dai_dynamic_2021, we can employ the perturbation and maximum principle to establish a necessary condition to characterize an open-loop time-consistent equilibrium strategy in Theorem 3.1 for our utility maximization problem with variance minization via a two-dimensional system of coupled BSDEs (3.1) (see also (4.2) for a simplified yet still coupled system), which has not been addressed before in the existing literature.
In general, the system (4.2) consists of two quadratic BSDEs coupled by the derived optimal strategy in quadratic form. Classical existence results for the multidimensional BSDE mainly fall into two categories: the fixed-point approach based on contraction mappings, which ensures uniqueness of solutions (see HU20161066, Luo_2017 and FAN2023105), and the construction of uniformly convergent approximations via coefficient regularization (see xing_class_2018, Jackson_2022 and JACKSON2023). However, neither of these methods applies to our problem (see Appendix C for the detailed explanation).
In response to the challenge of the unconventional BSDE system, we first study two special cases where the existence of solution can be obtained and the verification theorem can be exercised: (i) the case with zero correlation coefficient ($\rho=0$) for two Brownian motions $B$ and $\bar{B}$; (ii) the case with trading constraint in a bounded set.
For the general case with $\rho\neq 0$ and without trading constraint, we can no longer establish the existence of solution to the more complicated BSDE system. However, as a new contribution to the existing literature, we can show that by employing the solution structure from the time-consistent equilibrium strategy in the special case with $\rho=0$ we can successfully construct an approximate time-consistent equilibrium (see Definition 5.1) in the general case with $\rho\neq 0$ when the correlation coefficient $\rho$ is sufficiently small. More importantly, under some additional regularity conditions, we can rigorously show that the approximation error using the equilibrium solution in the case $\rho=0$ is of order $O(\rho^{2})$, see Theorem 5.5. This approximation result using the explicit equilibrium structure in the special case with $\rho=0$ significantly simplifies the decision making in some general incomplete market models with small $\rho$ such that it is sufficient to implement the more tractable control process while the time-consistency can be retained in an approximate sense.
We also highlight the main differences between the present paper and some related studies in Gu_2020 and Bensoussan_2025. Gu_2020 studies a time-consistent Hamilton–Jacobi–Bellman (HJB) equation to characterize the equilibrium strategy for a utility–deviation risk control problem. They obtain the explicit equilibrium value function and the corresponding equilibrium strategy in a homogeneous utility–deviation risk problem with CRRA utility based on a specific choice of risk-aversion function. In contrast, we focus on the optimization of the trade-off between the exponential utility and variance over log returns and characterize the open-loop time-consistent equilibrium via a system of coupled quadratic BSDEs, for which we establish existence results in some cases and provide an approximate equilibrium in the general case. Bensoussan_2025 study a stochastic control problem incorporating higher-order moments in the complete market model and provide the characterization of the time-consistent equilibrium strategy. On the contrary, we consider an incomplete stochastic factor model that significantly increases the mathematical complexity to establish the existence of time-consistent equilibrium.
To illustrate our theoretical results, we further employ a deep learning–based numerical scheme, originally proposed by e_deep_2017 and Han_2018, to solve the BSDE systems arising in the three cases—namely, the case with $\rho=0$, the case with trading constraints, and the case of the approximate equilibrium for $\rho\neq 0$. We find that when the correlation coefficient $\rho=0$, the equilibrium investment ratio exhibits a slight upward trend as time approaches the terminal horizon, differing from the observation in dai_dynamic_2021 such that the investment strategy remains constant. When $\rho<0$, however, our results are consistent with dai_dynamic_2021, showing a downward trend in the investment ratio when time approaches the terminal horizon. In addition, we perform sensitivity analysis with respect to the risk-aversion parameter $\zeta$ in the utility function and the weighting coefficient $\gamma$ between the expected utility and the variance. The numerical results illustrate that the equilibrium strategy exhibits a reasonable monotonic pattern with respect to both $\zeta$ and $\gamma$. Moreover, the numerical results also shows that the approximate Nash equilibrium strategy and its resulting value function are relatively close to those under the true constrained equilibrium when the trading constraint bound is large.
The remainder of the paper is organized as follows. Section 2 formulates the time inconsistent control problem as the trade-off between the utility maximization and the variance minimization over logarithmic portfolio returns, for which a time-consistent equilibrium is introduced. Section 3 provides the characterization of the time-consistent equilibrium by a quadratic BSDE system. Section 4 investigates the existence of solution to the BSDE system in the stochastic factor model in two special cases: (i) when two Brownian motions in the price dynamics and in the factor model are independent; (2) when the trading strategy is constrained in a compact set. Section 5 establishes the approximate time-consistent equilibrium in the general case when two Brownian motions are correlated with $\rho\neq 0$ using the same solution structure from time-consistent equilibrium solution in the model with $\rho=0$ with the error of order $O(\rho^{2})$. Some numerical illustrations and financial implications are presented in Section 6. Finally, Appendix A and B collect some inequalities and sufficient conditions for the existence theory of BSDE system that are used in previous sections, and Appendix C elaborates some challenges of the BSDE system in the general case when $\rho\neq 0$.
Notations: For $t\in[0,T]$, $p,q\geq 1$, we list below some notations that will be used frequently throughout the paper:
  • $H_{\textup{BMO}}$ denotes the set of $\mathbb{R}$-valued $\mathbb{F}$-progressively measurable processes such that $\|\pi\|_{\textup{BMO}}<\infty$, where
    $\|\pi\|_{\textup{BMO}}^{2}:=\sup_{\tau\in\mathcal{T}_{[0,T]}}\left\|\mathbb{E}_{\tau}\left[\int_{\tau}^{T}|\pi_{s}|^{2}ds\right]\right\|_{\infty},$
    $\mathcal{T}_{[0,T]}$ denotes all stopping times on $[0,T]$.
  • $L^{\infty}_{\mathcal{F}_{t}}(\Omega;\mathbb{R})$ denotes the set of all $\mathbb{R}$-valued $\mathcal{F}_{t}$-measurable bounded random variables.
  • $L^{p}_{\mathbb{F}}(\Omega;L^{q}(t,T;\mathbb{R}))$ denotes the set of all $\mathbb{R}$-valued progressively measurable processes $X$ with
    $\mathbb{E}\left[\left(\int_{t}^{T}|X_{s}|^{q}ds\right)^{p/q}\right]<\infty.$
    When $p=q$, we write $L^{p}_{\mathbb{F}}(t,T;\mathbb{R}):=L^{p}_{\mathbb{F}}(\Omega;L^{p}(t,T;\mathbb{R}))$
  • $L^{p}_{\mathbb{F}}(\Omega;C([t,T];\mathbb{R}))$ denotes the set of all $\mathbb{R}$-valued adapted and continuous processes $X$ with
    $\mathbb{E}\left[\sup_{s\in[t,T]}|X_{s}|^{p}\right]<\infty.$
  • $\mathcal{S}^{\infty}$ denotes the set of all $\mathbb{R}$-valued $\mathbb{F}$-progressively measurable and bounded processes.
2 Problem Formulation
Let $T>0$ be a finite time horizon and $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered complete probability space, where $\mathbb{F}=\{\mathcal{F}_{t},0\leq t\leq T\}$ is the augmented natural filtration generated by two standard Brownian motions $\{B(t),0\leq t\leq T\}$ and $\{\bar{B}(t),0\leq t\leq T\}$. Their correlation coefficient $\rho$ is such that $\mathbb{E}[B(t)\bar{B}(t)]=\rho t$. The market model consists of one risk-free bond with interest rate $r=\{r_{t}:t\in[0,T]\}$ and one stockwith price dynamics governed by
$\mathrm{d}S_{t}=S_{t}\left(\mu_{t}\mathrm{d}t+\sigma_{t}\mathrm{d}B_{t}\right),\quad S_{0}=s_{0}\in(0,\infty),$
where the drift $\mu:[0,T]\times\Omega\to\mathbb{R}$, the volatility $\mathbf{\sigma}>0:[0,T]\times\Omega\to\mathbb{R}$, and the interest rate $r:[0,T]\times\Omega\to\mathbb{R}$, are all $\mathbb{F}$-progressively measurable processes.
Assumption 2.1.
$r_{t},\mu_{t},\sigma_{t}>0$ and $\theta_{t}:=\sigma_{t}^{-1}(\mu_{t}-r_{t})$ are bounded $\mathbb{F}$-progressively measurable processes.
A trading strategy is a $\mathbb{F}$-progressively measurable processes $\pi$ such that $\int_{0}^{T}|\pi_{t}(\mu(t)-r(t))|\mathrm{d}t+\int_{0}^{T}|\sigma(t)\pi_{t}|^{2}\mathrm{d}t<\infty$ a.s., where $\pi_{t}$ stands for the fraction of wealth allocated in the stock at time $t$. The self-financing wealth process $W^{\pi}$ evolves according to
$\mathrm{d}W^{\pi}_{t}=W^{\pi}_{t}\left(r_{t}+\pi_{t}(\mu_{t}-r_{t})\mathrm{d}t+\pi_{t}\sigma_{t}\mathrm{d}B_{t}\right),\quad W^{\pi}_{0}=w_{0}>0.$
In this paper, we consider a utility maximization framework with variance adjustment on log return of the portfolio. That is, with given $r_{0}\in\mathbb{R},\pi_{0}$ and $t\in[0,T)$, the objective functional of the investor given the trading strategy $\pi$ at time $t$ is defined by
$J(t,\pi)=\mathbb{E}_{t}\left[U\left(\log\left(\frac{W^{\pi}_{T}}{W^{\pi}_{t}}\right)\right)\right]-\frac{\gamma_{t}}{2}\textup{Var}_{t}\left[\log\left(\frac{W^{\pi}_{T}}{W^{\pi}_{t}}\right)\right],$ (2.1)
where $\mathbb{E}_{t}$ and $\textup{Var}_{t}$ represent the conditional expectation and the conditional variance at time $t$, respectively, and $\gamma_{t}>0$ can be regarded as the weighting parameter measuring the trade-off between the expected utility and the risk (variance) at time $t$. In particular, we shall work with the CARA utility $U(x)=-\frac{1}{\zeta}e^{-\zeta x}$ with the risk aversion parameter $\zeta>0$.
To simplify the notation, let us denote $R^{\pi}_{t}=\log W^{\pi}_{t}$, which satisfies
$\mathrm{d}R^{\pi}_{t}=\left[r_{t}+\pi_{t}(\mu_{t}-r_{t})-\frac{1}{2}|\sigma_{t}\pi_{t}|^{2}\right]\mathrm{d}t+\pi_{t}\sigma_{t}\mathrm{d}B_{t},\quad R^{\pi}_{0}=r_{0}=\log w_{0}\in\mathbb{R}.$
Then the objective functional in (2.1) at time $t$ can be written as
$\mathbb{E}_{t}\left[U(R^{\pi}_{T}-R^{\pi}_{t})\right]-\frac{\gamma_{t}}{2}\textup{Var}_{t}[R^{\pi}_{T}-R^{\pi}_{t}]=-\frac{1}{\zeta}\mathbb{E}_{t}\left[\exp\left(-\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]-\frac{\gamma_{t}}{2}\textup{Var}_{t}[R^{\pi}_{T}].$
To ensure the objective functional under $\pi$ is well defined, we impose the following condition on trading strategies:
Definition 2.2.
(Admissible trading strategy) A trading strategy $\pi$ is admissible if $\pi\in H_{\textup{BMO}}$ and, for some $p>1$, it satisfies
$\mathbb{E}_{t}\left[\exp\left(-p\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]<\infty,\quad a.s.$ (2.2)
for any $t\in[0,T)$. We denote by $\Pi_{0}$ the set of all admissible trading strategies.
Remark 2.3.
Note that we have to require $\mathbb{E}_{t}\left[\exp\left(-\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]<\infty,a.s.$ to ensure that the objetive functional at time $t$ is well-defined. In (2.2), we impose a slightly stronger integrability condition (with $p>1$). As will be shown in the subsequent discussion, this condition guarantees that the admissible set remains open under perturbations (see Lemma 2.7 and Corollary 2.9), which is essential for the rigorous definition of the equilibrium. Moreover, it also implies that $\mathbb{E}_{t}\left[\left(U^{\prime}(R^{\pi}_{T}-R^{\pi}_{t})\right)^{p}\right]<\infty,a.s.$, a fact that will play a crucial role in the proof of Theorem 3.1; see Hamaguchi_2021 for similar conditions.
The following lemma shows that $J(t,\pi)<\infty$ for any $\pi\in\Pi_{0}$.
Lemma 2.4.
For $\pi\in\Pi_{0}$ and $t\in[0,T)$, $J(t,\pi)$ is finite almost surely. Moreover, $R^{\pi}\in L^{2}_{\mathbb{F}}(\Omega;C([0,T];\mathbb{R}))$.
Proof.
First, by Hölder’s inequality, we have
$\mathbb{E}_{t}\left[\exp\left(-\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]\leq\left(\mathbb{E}_{t}\left[\exp\left(-p\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]\right)^{\frac{1}{p}}<\infty,\quad a.s..$
By the definition of the admissible strategy $\pi$, it holds that
$R^{\pi}_{T}-R^{\pi}_{t}=\int_{t}^{T}\left[r_{s}+\pi_{s}(\mu_{s}-r_{s})-\frac{1}{2}|\sigma_{s}\pi_{s}|^{2}\right]\mathrm{d}s+\pi_{s}\sigma_{s}\mathrm{d}B_{s}.$
Using the standard estimate for SDE (see Theorem 3.4.3 in zhang_backward_2017), we obtain that
$\mathbb{E}\left[\sup\limits_{s\in[t,T]}\left|R^{\pi}_{s}-R^{\pi}_{t}\right|^{2}\mathds{1}_{A}\right]=$ $C(T)\mathbb{E}\left[\left(\int_{t}^{T}\left[r_{s}+\pi_{s}(\mu_{s}-r_{s})-\frac{1}{2}|\sigma_{s}\pi_{s}|^{2}\right]\mathds{1}_{A}\mathrm{d}s\right)^{2}+\int_{t}^{T}|\pi_{s}\sigma_{s}|^{2}\mathds{1}_{A}\mathrm{d}s\right]$
$\leq$ $C(\|r\|_{\infty},\|\mu\|_{\infty},\|\sigma\|_{\infty},T)\mathbb{E}\left[\left(\left(\int_{t}^{T}|\pi_{s}|^{2}\mathrm{d}s\right)^{2}+1\right)\mathds{1}_{A}\right].$
Note that $A\in\mathcal{F}_{t}$ is arbitrary and $\pi\in H_{\textup{BMO}}$, by the definition of conditional expectation and Lemma A.1, we deduce that
$\left(\mathbb{E}_{t}\left[\left|R^{\pi}_{T}-R^{\pi}_{t}\right|\right]\right)^{2}\leq\mathbb{E}_{t}\left[\left|R^{\pi}_{T}-R^{\pi}_{t}\right|^{2}\right]$ $\leq\mathbb{E}_{t}\left[\sup\limits_{s\in[t,T]}\left|R^{\pi}_{s}-R^{\pi}_{t}\right|^{2}\right]$
$\leq C(\|r\|_{\infty},\|\mu\|_{\infty},\|\sigma\|_{\infty},T)\left(\mathbb{E}_{t}\left[\left(\int_{t}^{T}|\pi_{s}|^{2}\mathrm{d}s\right)^{2}\right]+1\right)$
$\leq C(\|r\|_{\infty},\|\mu\|_{\infty},\|\sigma\|_{\infty},T)\left(\|\pi\|_{\textup{BMO}}^{4}+1\right)<\infty.$
Thus $\textup{Var}_{t}[R^{\pi}_{T}-R^{\pi}_{t}]$ is bounded, and consequently, $J(t,\pi)$ is finite for any $t\in[0,T)$. The proof also verifies that $R^{\pi}\in L^{2}_{\mathbb{F}}(\Omega;C([0,T];\mathbb{R}))$. ∎
Remark 2.5.
Note that for a given admissible trading strategy $\pi$, $R^{\pi}_{T}-R^{\pi}_{t}$ is independent of the initial value $r_{0}$. Hence, the definition of the admissible set $\Pi_{0}$ is independent of the initial value $r_{0}$.
The issue of time inconsistency arises due to the initial dependence of $R_{t}$ and the variance term, as a result, the optimal solution make no longer sense as the optimal decision today might be turned over at future dates. Instead, we seek the equilibrium solution using the game theoretic thinking with future selves. To define the equilibrium strategy, we first introduce the perturbation strategy of an admissible strategy.
Definition 2.6.
Assume that $\pi$ is an admissible strategy, for any fixed $t\in[0,T)$, $\varepsilon\in(0,T-t)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$, define an admissible perturbation strategy $\pi^{t,\varepsilon,\eta}$ (see Corollary 2.9) as follows:
$\pi_{s}^{t,\varepsilon,\eta}={\pi}_{s}+\eta\mathds{1}_{s\in[t,t+\varepsilon)},\quad s\in[0,T],$
where $\mathds{1}$ denotes the indicator function.
Let $\xi^{t,\varepsilon,\eta,\pi}:=R^{\pi^{t.\varepsilon,\eta}}-R^{\pi}$ be the difference of log return processes between $\pi^{t,\varepsilon,\eta}$ and $\pi$. Then for any $t\in[0,T)$, $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R}^{d})$, $\varepsilon\in(0,T-t)$ and an admissible strategy $\pi$, $\xi^{t,\varepsilon,\pi}:=\xi^{t,\varepsilon,\eta,\pi}$ [^3] For notation simplicity, we drop the superscript $\eta$. is the unique solution of the SDE [^4] Formally, both $R^{\pi^{t.\varepsilon,\eta}}$ and $\xi^{t,\varepsilon,\pi}$ can be defined on the entire interval $[0,T]$. Nevertheless, in view of the definition of equilibrium, it suffices to consider them only on $[t,T]$. that
$\left\{\begin{aligned} &\mathrm{d}\xi_{s}^{t,\varepsilon,\pi}=\left(a(s,\pi^{t,\varepsilon,\eta}_{s})-a(s,\pi_{s})\right)\mathrm{d}s+\sigma_{s}\eta 1_{s\in[t,t+\varepsilon)}\mathrm{d}B_{s},\\ &\xi_{t}^{t,\varepsilon,\pi}=0.\end{aligned}\right$. (2.3)
Here, $a(s,\pi):=r_{s}+(\mu_{s}-r_{s})\pi-\frac{1}{2}|\sigma_{s}\pi|^{2}$ . In the following lemma, we will first establish some fundamental properties of $\xi^{t,\varepsilon,\pi}$.
Lemma 2.7.
Let $t\in[0,T)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$ be fixed.
  1. 1.
    For any $\varepsilon\in(0,T-t)$, $\xi^{t,\varepsilon,\pi}\in L_{\mathbb{F}}^{k}(\Omega;C([t,T];\mathbb{R}))$ for any $k\geq 1$.
  2. 2.
    For any $k\geq 1$, there exists a constant $C(\|r\|_{\infty},\|\mu\|_{\infty},\|\sigma\|_{\infty},T,k,\|\pi\|_{\textup{BMO}})$ such that E_t[sup_s[t,T]—ξ_s^t,ε,π—^2k]C(r_,μ_,σ_,T,k,π_BMO)(ε—η—^2)^k, a.s. for $\varepsilon\in(0,T-t)$ sufficiently small.
  3. 3.
    For any $c>0,\pi\in\Pi_{0},\varepsilon\in(0,T-t)$, it holds that $\sup\limits_{\varepsilon\in(0,T-t)}\mathbb{E}_{t}\left[\exp(c|\xi^{t,\varepsilon,\pi}_{T}|)\right]<C(c,\pi,\|\eta\|_{\infty})<\infty,\mathbb{P}-a.s.$, where $C(c,\pi,\|\eta\|_{\infty})$ denotes a constant depending only on $c,\pi,\|\eta\|_{\infty}$, but independent of $\epsilon$, and is increasing with respect to $\|\eta\|_{\infty}$.
Proof.
The first assertion is standard, see Theorem 3.4.3 in zhang_backward_2017. The proof of the second assertion is similar to the ones of Lemma 2.4 and Lemma 2.5 in Hamaguchi_2021. Indeed, for any set $A\in\mathcal{F}_{t}$, it holds that
$\mathbb{E}\left[\sup\limits_{s\in[t,T]}|\xi_{s}^{t,\varepsilon,\pi}|^{2k}\mathds{1}_{A}\right]\leq$ $C\mathbb{E}\left[\left(\int_{t}^{t+\varepsilon}\left|(\mu_{s}-r_{s})\eta-\frac{1}{2}\sigma_{s}^{2}\eta^{2}-\sigma_{s}^{2}\pi_{s}\eta\right|\mathds{1}_{A}\mathrm{d}s\right)^{2k}+\left(\int_{t}^{t+\varepsilon}|\sigma_{s}\eta|^{2}\mathds{1}_{A}\mathrm{d}s\right)^{k}\right]$
$\leq$ $C\mathbb{E}\left[\left\{\left(\varepsilon+\varepsilon\eta+\varepsilon^{\frac{1}{2}}\left(\int_{t}^{t+\varepsilon}|\pi|^{2}\mathrm{d}s\right)^{\frac{1}{2}}\right)^{2k}|\eta|^{2k}+\varepsilon^{k}|\eta|^{2k}\right\}\mathds{1}_{A}\right]$
$=$ $C\mathbb{E}\left[\left((\varepsilon|\eta|^{2})^{k}\left[\left(\int_{t}^{t+\varepsilon}|\pi_{s}|^{2}\mathrm{d}s\right)^{k}+1\right]+o(\varepsilon^{k})\right)\mathds{1}_{A}\right].$
Here, we have omitted the dependence of $C$ for simplicity. Then, it follows that
$\mathbb{E}_{t}\left[\sup\limits_{s\in[t,T]}|\xi_{s}^{t,\varepsilon,\pi}|^{2k}\right]\leq C(\varepsilon|\eta|^{2})^{k}\left(1+\mathbb{E}_{t}\left[\left(\int_{t}^{T}|\pi_{s}|^{2}\mathrm{d}s\right)^{k}\right]\right)\leq C(\varepsilon|\eta|^{2})^{k},$
where we have used the fact that $\pi\in H_{\textup{BMO}}$ and the conditional version of the energy inequality (see Lemma A.1).
For the third assertion, note that
$\mathbb{E}_{t}\left[\exp(c|\xi^{t,\varepsilon,\pi}_{T}|)\right]$
$=$ $\mathbb{E}_{t}\left[\exp\left(c\int_{t}^{t+\varepsilon}\left((\mu_{s}-r_{s})\eta-\frac{1}{2}\sigma_{s}^{2}\eta^{2}-\sigma_{s}^{2}\pi_{s}\eta\right)\mathrm{d}s+c\int_{t}^{t+\varepsilon}\sigma_{s}\eta\mathrm{d}B_{s}\right)\right]$
$\leq$ $C(c,\eta)\left(\mathbb{E}_{t}\left[\exp\left(2c\|\sigma\|^{2}_{\infty}|\eta|\int_{t}^{T}|\pi_{s}|\mathrm{d}s\right)\right]\right)^{\frac{1}{2}}\left(\mathbb{E}_{t}\left[\exp\left(2c\int_{t}^{t+\varepsilon}\sigma_{s}\eta\mathrm{d}B_{s}-2c^{2}\int_{t}^{t+\varepsilon}|\sigma_{s}\eta|^{2}\right)\right]\right)^{\frac{1}{2}}$
$\leq$ $C(c,\delta,\|\eta\|_{\infty})\left(\mathbb{E}_{t}\left[\exp\left(\delta\int_{t}^{T}|\pi_{s}|^{2}\mathrm{d}s\right)\right]\right)^{\frac{1}{2}}\left(\mathbb{E}_{t}\left[\exp\left(2c\int_{t}^{t+\varepsilon}\sigma_{s}\eta\mathrm{d}B_{s}-2c^{2}\int_{t}^{t+\varepsilon}|\sigma_{s}\eta|^{2}\mathrm{d}s\right)\right]\right)^{\frac{1}{2}}.$
In view that $\sigma$ and $\eta$ are bounded, the Novikov condition is satisfied such that the second term equals $1$ a.s.. We also note that $C(c,\delta)$ can be chosen independent of $\varepsilon$. In fact, we can choose
$C(c,\delta,\|\eta\|_{\infty})=\exp\left(c(T-t)\left(\|\mu-r\|_{\infty}\cdot\|\eta\|_{\infty}+\frac{1}{2}\|\sigma\|_{\infty}^{2}\|\eta\|_{\infty}^{2}\right)+c^{2}(T-t)\|\sigma\|_{\infty}^{2}\|\eta\|_{\infty}^{2}+\frac{T}{\delta}\left(2c\|\sigma\|^{2}_{\infty}\|\eta\|_{\infty}\right)^{2}\right).$
By the John–Nirenberg inequality (see Lemma A.2), for $\delta>0$ such that $\delta\|\pi\|_{\textup{BMO}}^{2}<1$, we have
$\mathbb{E}_{t}\left[\exp\left(\delta\int_{t}^{T}|\pi_{s}|^{2}\mathrm{d}s\right)\right]<\frac{1}{1-\delta\|\pi\|_{\textup{BMO}}^{2}}<\infty\quad a.s..$
As a result,
$\sup\limits_{\varepsilon\in(0,T-t)}\mathbb{E}_{t}\left[\exp(c|\xi^{t,\varepsilon,\pi}_{T}|)\right]\leq C(c,\delta,\|\eta\|_{\infty})\left(\frac{1}{1-\delta\|\pi\|_{\textup{BMO}}^{2}}\right)^{\frac{1}{2}}<\infty\quad a.s..$
Remark 2.8.
The above result is similar to Lemma 2.5 in Hamaguchi_2021. However, as the control process appears in the SDE governing the difference process, $\xi^{t,\varepsilon,\pi}$ cannot satisfy the properties in Hamaguchi_2021 under arbitrary trading strategies. This motivates us to consider the BMO condition into the definition of our admissible strategies. Consequently, the proof differs from that of Lemma 2.5 in Hamaguchi_2021.
Corollary 2.9.
Assume that $\pi$ is an admissible trading strategy, for any $t\in[0,T)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$, $\pi^{t,\varepsilon,\eta}$ is an admissible strategy.
Proof.
We can choose $1<p^{\prime}<p$ and $q>1$ such that $\frac{1}{p}+\frac{1}{q}=\frac{1}{p^{\prime}}$. By Lemma 2.7, it holds that
$\left(\mathbb{E}_{t}\left[\exp\left(-p^{\prime}\zeta(R^{\pi^{t,\varepsilon,\eta}}_{T}-R^{\pi^{t,\varepsilon,\eta}}_{t})\right)\right]\right)^{\frac{1}{p^{\prime}}}$
$=$ $\left(\mathbb{E}_{t}\left[\exp\left(-p^{\prime}\zeta\left[(R^{\pi}_{T}-R^{\pi}_{t})+\xi_{T}^{t,\varepsilon,\pi}\right]\right)\right]\right)^{\frac{1}{p^{\prime}}}$
$\leq$ $\left(\mathbb{E}_{t}\left[\exp\left(-p\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]\right)^{\frac{1}{p}}\left(\mathbb{E}_{t}\left[\exp\left(q\zeta\left|\xi_{T}^{t,\varepsilon,\pi}\right|\right)\right]\right)^{\frac{1}{q}}<\infty\;a.s..$
Moreover, it is obvious that under a bounded perturbation $\eta$,
$\|\pi^{t,\varepsilon,\eta}\|_{\textup{BMO}}^{2}\leq 2\sup_{\tau\in\mathcal{T}_{[0,T]}}\left\|\mathbb{E}_{\tau}\left[\int_{\tau}^{T}|\pi_{s}|^{2}ds\right]\right\|_{\infty}+2\varepsilon\|\eta\|_{\infty}^{2}<\infty.$
Thus, $\pi^{t,\varepsilon,\eta}$ is an admissible strategy. ∎
Next we give the definition of our (open-loop) time-consistent equilibrium strategy using the intra-personal game theoretic thinking by perturbation, which is inspired by hu_time-inconsistent_2012; hu_time-inconsistent_2017.
Definition 2.10.
For $\hat{\pi}\in\Pi_{0}$, $\hat{\pi}$ is called a time-consistent equilibrium strategy if
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{J(t,\hat{\pi}^{t,\varepsilon,\eta})-J(t,\hat{\pi})}{\varepsilon}\leq 0\quad a.s.,$
for any $t\in[0,T)$ and any $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$.
Remark 2.11.
In the following sections, we often fix a candidate equilibrium strategy $\hat{\pi}$ or an arbitrary admissible strategy $\pi$ for analysis. In proving the necessary conditions or verifying equilibrium, we will work with the perturbed log return process under a fixed perturbation $\eta$. To ease presentation, we denote by $\hat{\pi}^{t,\epsilon},\pi^{t,\epsilon}$ the perturbation strategy $\hat{\pi}^{t,\epsilon,\eta},\pi^{t,\epsilon,\eta}$, and denote by $\hat{R},R$ the log return process ${R}^{\hat{\pi}},R^{\pi}$, and denote by $\hat{R}^{t,\varepsilon},R^{t,\varepsilon}$ the perturbed log return process ${R}^{\hat{\pi}^{t,\epsilon,\eta}},{R}^{\pi^{t,\epsilon,\eta}}$, and denote by $\xi^{t,\varepsilon}$ the difference process $\xi^{t,\varepsilon,\pi}$ under the perturbation. For simplicity, we also use the notations $\hat{u}:=\sigma\hat{\pi}$ and $u:=\sigma\pi$ in the following sections.
3 A Necessary Condition of Time-Consistent Equilibria
In this section, we establish several properties of admissible strategies and characterize the necessary conditions of a time-consistent equilibrium by a BSDE system. Without loss of generality, we consider a fixed $p$ such that any admissible strategy satisfies
$\mathbb{E}_{t}\left[\exp\left(-p\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]<\infty,\quad a.s..$
Theorem 3.1.
[…]
  1. 1.
[…]
  2. 2.
[…]
    $\left\{\begin{aligned} &\mathrm{d}Y_{s}=\left(\frac{\zeta}{2}|Z_{s}^{1}+\sigma_{s}\hat{\pi}_{s}+\rho Z_{s}|^{2}+\frac{\zeta(1-\rho^{2})}{2}|Z_{s}|^{2}-a(s,\hat{\pi}_{s})\right)\mathrm{d}s+Z_{s}^{1}\mathrm{d}B_{s}+Z_{s}\mathrm{d}\bar{B}_{s},\\ &\mathrm{d}\tilde{Y}_{s}=-a(s,\hat{\pi}_{s})\mathrm{d}s+\tilde{Z}^{1}_{s}\mathrm{d}B_{s}+\tilde{Z}_{s}\mathrm{d}\bar{B}_{s},\\ &Y_{T}=0,\\ &\tilde{Y_{T}}=0.\end{aligned}\right$. (3.1)
Furthermore, the equilibrium strategy is characterized by
$\hat{\pi}=\sigma^{-1}\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}(Z^{1}+\rho Z)-\gamma(\tilde{Z}^{1}+\rho\tilde{Z})}{(\zeta+1)e^{-\zeta Y}+\gamma}.$
Remark 3.2.
When the market model becomes complete, i.e., the second Brownian motion $\bar{B}$ is absent, the BSDE system (3.1) degenerates to
$\left\{\begin{aligned} &\mathrm{d}Y_{s}=\left(\frac{\zeta}{2}|Z_{s}+\hat{u}_{s}|^{2}-a(s,\hat{\pi}_{s})\right)\mathrm{d}s+Z_{s}\mathrm{d}B_{s},\\ &\mathrm{d}\tilde{Y}_{s}=-a(s,\hat{\pi}_{s})\mathrm{d}s+\tilde{Z}\mathrm{d}B_{s},\\ &Y_{T}=0,\\ &\tilde{Y_{T}}=0.\end{aligned}\right$. (3.2)
The equilibrium strategy takes the form $\hat{\pi}=\sigma^{-1}\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}Z-\gamma\tilde{Z}}{(\zeta+1)e^{-\zeta Y}+\gamma}$. We observe that the BSDE system admits a solution when $Z=0$ and $\tilde{Z}=0$, thus an equilibrium is given by
$\hat{\pi}=\sigma^{-1}\frac{e^{-\zeta Y}\theta}{(\zeta+1)e^{-\zeta Y}+\gamma},$
where $Y$ is the solution of the ODE with terminal condition:
$\left\{\begin{aligned} &\mathrm{d}Y_{s}=\left(\frac{\zeta}{2}|\hat{u}_{s}|^{2}-a(s,\hat{\pi}_{s})\right)\mathrm{d}s,\\ &Y_{T}=0.\end{aligned}\right$. (3.3)
Let $A=e^{-\zeta Y}$, then $A$ satisfies
$A_{t}=A(f(A)):=-\frac{\zeta(\zeta+1)A^{3}\theta^{2}}{2((\zeta+1)A+\gamma)^{2}}+\frac{\zeta A^{2}\theta^{2}}{(\zeta+1)A+\gamma}+r\zeta A,$ (3.4)
with the terminal condition $A(T)=1$. It is clear that $A$ has a lower bound 0. Combining $A^{\prime}=A(f(A))$ and the fact that $f(A)$ is bounded, we deduce that $A$ is bounded from above and thus there exists a unique solution and $A$ is positive.
Remark 3.3.
Hamaguchi_2021 derives a one-dimensional FBSDE as a necessary condition for an equilibrium in a general non-exponential time-inconsistent problem. In our setting, however, the introduced variance must to be represented by another BSDE of $\tilde{Y}$, which leads to strong coupling in our equation. Although dai_dynamic_2021 study a mean-variance framework including a variance term in their objective function, it is straightforward to see that the mean term can be represented by the BSDE of $\tilde{Y}$. In fact, $\tilde{Y}:=\mathbb{E}_{t}[R_{T}-R_{t}]$ (see in Lemma 3.4 later) corresponds precisely to the mean term at time $t$. Hence, their analysis also focus on a one-dimensional BSDE. Unlike Hamaguchi_2021 and dai_dynamic_2021, we focus on a two-dimensional fully coupled quadratic BSDE system.
To prove Theorem 3.1, we need some auxiliary preparations.
Lemma 3.4.
[…]
$J(t,R_{t},\pi^{t,\varepsilon})-J(t,R_{t},\pi)$
$=$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta(R_{s}+Y_{s}-R_{t})}(a(s,\pi^{t,\varepsilon}_{s})-a(s,\pi_{s})-\zeta(Z_{s}^{1}+\sigma_{s}\pi_{s}+\rho Z)\cdot\sigma_{s}\eta)\mathrm{d}s\right]$
$-\zeta\mathbb{E}_{t}\left[\int_{0}^{1}e^{-\zeta(R_{T}+\lambda\xi_{T}^{t,\varepsilon}-R_{t})}(1-\lambda)d\lambda|\xi_{T}^{t,\varepsilon}|^{2}\right]$
$+\frac{\gamma}{2}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}((\sigma_{s}\pi_{s}+\tilde{Z}_{s}^{1}+\rho\tilde{Z}_{s})^{2}-(\sigma_{s}\pi_{s}^{t,\varepsilon}+\tilde{Z}_{s}^{1}+\rho\tilde{Z}_{s})^{2})\mathrm{d}s\right].$
Proof.
Fix the strategy $\pi$, $t\in[0,T)$, $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R}^{d})$ and $\varepsilon\in(0,T-t)$. By the definition of $J$, we have
$J(t,R_{t},\pi^{t,\varepsilon})-J(t,R_{t},\pi)=\mathbb{E}_{t}\left[U(R_{T}^{t,\varepsilon}-R_{t})-U(R_{T}-R_{t})\right]-\frac{\gamma}{2}\left(\textup{Var}_{t}[R_{T}^{t,\varepsilon}]-\textup{Var}_{t}[R_{T}]\right)$
For the first term, we have that
$\mathbb{E}_{t}\left[U(R_{T}^{t,\varepsilon}-R_{t})-U(R_{T}-R_{t})\right]=e^{\zeta R_{t}}\mathbb{E}_{t}\left[U^{{}^{\prime}}(R_{T})\xi_{T}^{t,\varepsilon}+\int_{0}^{1}U^{{}^{\prime\prime}}(R_{T}+\lambda\xi_{T}^{t,\varepsilon})(1-\lambda)d\lambda|\xi_{T}^{t,\varepsilon}|^{2}\right].$
Next we analyze the conditional expectation $\mathbb{E}_{t}\left[U^{{}^{\prime}}(R_{T})\xi_{T}^{t,\varepsilon}\right]$ and let $(\alpha,\beta)$ be the unique adapted solution of the BSDE
$\left\{\begin{aligned} &\mathrm{d}\alpha_{s}=\beta_{s}^{1}\mathrm{d}B_{s}+\beta_{s}\mathrm{d}\bar{B}_{s},\\ &\alpha_{T}=e^{-\zeta R_{T}}.\end{aligned}\right$.
The martingale representation theorem impllies that $\alpha_{s}=\mathbb{E}_{s}\left[U^{{}^{\prime}}(R_{T})\right]$. By the definition of the admissible strategy, it holds that $\alpha_{T}\in L^{p}_{\mathcal{F}_{T}}(\Omega;\mathbb{R})$ for some $p>1$. Let $Y=-\frac{1}{\zeta}\log\alpha-R$ and hence $U^{{}^{\prime}}(R_{s}+Y_{s})=\alpha_{s}=\mathbb{E}_{s}\left[U^{{}^{\prime}}(R_{T})\right]$. In particular, $\alpha$ is positive and $Y$ is well defined for any $t\in[0,T]$ a.s.. Moreover, $\alpha\in L^{p}_{\mathbb{F}}(\Omega;C([0,T];\mathbb{R}))$ and $\beta^{1},\beta\in L^{p}_{\mathbb{F}}\left(\Omega;L^{2}(0,T;\mathbb{R})\right)$. Then $Y$ is an adapted process satisfying the following BSDE that
$\left\{\begin{aligned} &\mathrm{d}Y_{s}=\left(\frac{1}{2\zeta\alpha_{s}^{2}}\left(|\beta_{s}^{1}|^{2}+|\beta_{s}|^{2}+2\rho\beta_{s}^{1}\beta_{s}\right)-a(s,\pi_{s})\right)\mathrm{d}s-\left(\frac{\beta_{s}^{1}}{\zeta\alpha_{s}}+\sigma\pi_{s}\right)\mathrm{d}B_{s}-\frac{\beta_{s}}{\zeta\alpha_{s}}\mathrm{d}\bar{B}_{s},\\ &Y_{T}=0.\end{aligned}\right$.
Let $Z^{1}=-\frac{\beta^{1}}{\zeta\alpha}-\sigma\pi$ and $Z=-\frac{\beta}{\zeta\alpha}$, it then holds that
$\mathrm{d}Y_{s}=\left(\frac{\zeta}{2}|Z_{s}^{1}+\sigma_{s}\pi+\rho Z_{s}|^{2}+\frac{\zeta(1-\rho^{2})}{2}|Z_{s}|^{2}-a(s,\pi_{s})\right)\mathrm{d}s+Z_{s}^{1}\mathrm{d}B_{s}+Z^{1}\mathrm{d}\bar{B}_{s}.$
Moreover, by taking the conditional expectations, we see that
$\mathbb{E}_{t}\left[U^{{}^{\prime}}(R_{T})\xi_{T}^{t,\varepsilon}\right]=\mathbb{E}_{t}[\alpha_{T}\xi_{t+\varepsilon}^{t,\varepsilon}]=\mathbb{E}_{t}\left[\mathbb{E}[\alpha_{T}|\mathcal{F}_{t+\varepsilon}]\xi_{t+\varepsilon}^{t,\varepsilon}\right]=\mathbb{E}_{t}[\alpha_{t+\varepsilon}\xi_{t+\varepsilon}^{t,\varepsilon}]=\mathbb{E}_{t}\left[U^{{}^{\prime}}(R_{t+\varepsilon}+Y_{t+\varepsilon})\xi_{t+\varepsilon}^{t,\varepsilon}\right].$
[…]
$\mathbb{E}_{t}\left[U^{{}^{\prime}}(R_{t+\varepsilon}+Y_{t+\varepsilon})\xi_{t+\varepsilon}^{t,\varepsilon}\right]$
$=$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}U^{{}^{\prime}}(R_{s}+Y_{s})\left(\left(a(s,\pi^{t,\varepsilon}_{s})-a(s,\pi_{s})\right)\mathrm{d}s+\sigma_{s}\eta\mathrm{d}B_{s}\right)\right$.
$+\int_{t}^{t+\varepsilon}\xi_{s}^{t,\varepsilon}U^{{}^{\prime\prime}}(R_{s}+Y_{s})\left((Z_{s}^{1}+\sigma_{s}\pi_{s})\mathrm{d}B_{s}+Z_{s}\mathrm{d}\bar{B}_{s}+\left(\frac{\zeta}{2}|Z_{s}^{1}+\sigma_{s}\pi_{s}+\rho Z_{s}|^{2}+\frac{\zeta(1-\rho^{2})}{2}|Z_{s}|^{2}\right)\mathrm{d}s\right)$
$+\frac{1}{2}\int_{t}^{t+\varepsilon}\xi_{s}^{t,\varepsilon}U^{{}^{\prime\prime\prime}}(R_{s}+Y_{s})\left(|Z_{s}^{1}+\sigma_{s}\pi_{s}|^{2}+|Z_{s}|^{2}+2\rho(Z_{s}^{1}+\sigma_{s}\pi_{s})Z_{s}\right)\mathrm{d}s$
$\left.+\int_{t}^{t+\varepsilon}U^{{}^{\prime\prime}}(R_{s}+Y_{s})(Z_{s}^{1}+\sigma_{s}\pi_{s}+\rho Z_{s})\cdot\sigma_{s}\eta\mathrm{d}s\right]$
$=$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta(R_{s}+Y_{s})}\left(a(s,\pi^{t,\varepsilon}_{s})-a(s,\pi_{s})-\zeta(Z_{s}^{1}+\sigma_{s}\pi_{s}+\rho Z_{s})\cdot\sigma_{s}\eta\right)\mathrm{d}s\right]$
$+\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta(R_{s}+Y_{s})}\left[(\sigma_{s}\eta-\zeta\xi_{s}^{t,\varepsilon}(Z_{s}^{1}+\sigma_{s}\pi_{s}))\mathrm{d}B_{s}-\zeta\xi_{s}^{t,\varepsilon}Z_{s}\mathrm{d}\bar{B}_{s}\right]\right].$ (3.5)
Under Assumption 2.1, we have
$\mathbb{E}_{t}\left[\left(\int_{t}^{t+\varepsilon}(\alpha_{s}\sigma_{s}\eta)^{2}\mathrm{d}s\right)^{\frac{1}{2}}\right]\leq\left(\mathbb{E}_{t}\left[\left(\sup\limits_{s\in[t,t+\varepsilon]}\alpha_{s}\right)^{p}\right]\right)^{\frac{1}{p}}\left(\mathbb{E}_{t}\left[\left(\sup\limits_{s\in[t,t+\varepsilon]}\sigma_{s}\eta\right)^{q}\right]\right)^{\frac{1}{q}}<\infty,$
$\mathbb{E}_{t}\left[\left(\int_{t}^{t+\varepsilon}(\beta_{s}\xi_{s}^{t,\varepsilon})^{2}\mathrm{d}s\right)^{\frac{1}{2}}\right]\leq\left(\mathbb{E}_{t}\left[\left(\int_{t}^{t+\varepsilon}(\beta_{s})^{2}\mathrm{d}s\right)^{\frac{p}{2}}\right]\right)^{\frac{1}{p}}\left(\mathbb{E}_{t}\left[\left(\sup\limits_{s\in[t,t+\varepsilon]}\xi_{s}^{t,\varepsilon}\right)^{q}\right]\right)^{\frac{1}{q}}<\infty,$
$\mathbb{E}_{t}\left[\left(\int_{t}^{t+\varepsilon}(\beta_{s}^{1}\xi_{s}^{t,\varepsilon})^{2}\mathrm{d}s\right)^{\frac{1}{2}}\right]\leq\left(\mathbb{E}_{t}\left[\left(\int_{t}^{t+\varepsilon}(\beta_{s}^{1})^{2}\mathrm{d}s\right)^{\frac{p}{2}}\right]\right)^{\frac{1}{p}}\left(\mathbb{E}_{t}\left[\left(\sup\limits_{s\in[t,t+\varepsilon]}\xi_{s}^{t,\varepsilon}\right)^{q}\right]\right)^{\frac{1}{q}}<\infty.$
The Burkholder-Davis-Gundy inequality yields that
$\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta(R_{s}+Y_{s})}\left[(\sigma\eta-\zeta\xi_{s}^{t,\varepsilon}(Z_{s}^{1}+\sigma\pi))\mathrm{d}B_{s}-\zeta\xi_{s}^{t,\varepsilon}Z_{s}\mathrm{d}\bar{B}_{s}\right]\right]=0.$
For the second part, let $\tilde{\alpha}_{t}=\mathbb{E}_{t}[R_{T}]$ and $\tilde{Y_{t}}=\tilde{\alpha}_{t}-R_{t}$. Then, $\tilde{\alpha}$ and $\tilde{Y}$ satisfy that
$\mathrm{d}\tilde{\alpha}=\tilde{\beta}^{1}\mathrm{d}B_{s}+\tilde{\beta}\mathrm{d}\bar{B}_{s},$
$\mathrm{d}\tilde{Y}_{s}=-a(s,\pi_{s})\mathrm{d}s+\tilde{Z}^{1}_{s}\mathrm{d}B_{s}+\tilde{Z}_{s}\mathrm{d}\bar{B}_{s},$
and $\tilde{\alpha}\in L^{2}_{\mathbb{F}}(\Omega;C([0,T];\mathbb{R})),\tilde{\beta}^{1}=\tilde{Z}^{1}+\sigma\pi,\tilde{\beta}=\tilde{Z}\in L^{2}_{\mathbb{F}}\left(\Omega;L^{2}(0,T;\mathbb{R})\right)$. From Lemma 2.4, it follows that $R^{\pi}\in L^{2}_{\mathbb{F}}(\Omega;C([0,T];\mathbb{R}))$ and hence $\tilde{Y}\in L^{2}_{\mathbb{F}}(\Omega;C([0,T];\mathbb{R}))$. We thus get that
$\textup{Var}_{t}[R_{T}]-\textup{Var}_{t}[R_{T}^{t,\varepsilon}]$
$=$ $\textup{Var}_{t}\left[\mathbb{E}_{t+\varepsilon}[R_{T}]\right]+\mathbb{E}_{t}\left[\textup{Var}_{t+\varepsilon}[R_{T}]\right]-\textup{Var}_{t}\left[\mathbb{E}_{t+\varepsilon}[R_{T}^{t,\varepsilon}]\right]-\mathbb{E}_{t}\left[\textup{Var}_{t+\varepsilon}[R_{T}^{t,\varepsilon}]\right]$
$=$ $\textup{Var}_{t}[R_{t+\varepsilon}+\tilde{Y}_{t+\varepsilon}]-\textup{Var}_{t}[R_{t+\varepsilon}^{t,\varepsilon}+\tilde{Y}_{t+\varepsilon}]$
$=$ $\textup{Var}_{t}\left[\int_{t}^{t+\varepsilon}(\sigma\pi_{s}+\tilde{Z}^{1}_{s})\mathrm{d}B_{s}+\tilde{Z}_{s}\mathrm{d}\bar{B}_{s}\right]-\textup{Var}_{t}\left[\int_{t}^{t+\varepsilon}(\sigma\pi_{s}^{t,\varepsilon}+\tilde{Z}_{s}^{1})\mathrm{d}B_{s}+\tilde{Z}_{s}\mathrm{d}\bar{B}_{s}\right]$
$=$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}((\sigma\pi_{s}+\tilde{Z}^{1}_{s}+\rho\tilde{Z}_{s})^{2}-(\sigma\pi_{s}^{t,\varepsilon}+\tilde{Z}^{1}_{s}+\rho\tilde{Z}_{s})^{2})\mathrm{d}s\right],$ (3.6)
where we have used the fact that $R_{T}-R_{t+\varepsilon}=R_{T}^{t,\varepsilon}-R_{t+\varepsilon}^{t,\varepsilon}$ in the second and third equalities . The conclusion then follows by combining (3) and (3). ∎
Lemma 3.5.
Let $\pi\in\Pi_{0}$ fixed. For any $t\in[0,T)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$ fixed and for $\varepsilon$ sufficiently small, it holds that a.s.
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{0}^{1}e^{-\zeta(R_{T}+\lambda\xi_{T}^{t,\varepsilon}-R_{t})}(1-\lambda)d\lambda|\xi_{T}^{t,\varepsilon}|^{2}\right]\leq C(\|\eta\|_{\infty})|\eta|^{2},\quad a.s.$
with $C$ denoting a constant independent of $\varepsilon$ and increasing with respect to $\|\eta\|_{\infty}$.
Proof.
Hölder’s inequality and Lemma 2.7 imply that $\mathbb{E}_{t}\left[e^{-\zeta R_{T}}|\xi_{T}^{t,\varepsilon}|^{2}\right]<\infty$ and $\mathbb{E}_{t}\left[e^{-\zeta(R_{T}+\xi_{T}^{t,\varepsilon})}|\xi_{T}^{t,\varepsilon}|^{2}\right]<\infty$ a.s.. Noting that $e^{-(R_{T}+\lambda\xi_{T}^{t,\varepsilon})}|\xi_{T}^{t,\varepsilon}|^{2}$ is nonnegative and convex almost sure with respect to $\lambda$, we have
$\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{0}^{1}\left(e^{-\zeta(R_{T}+\lambda\xi_{T}^{t,\varepsilon}-R_{t})}(1-\lambda)\right)d\lambda\cdot|\xi_{T}^{t,\varepsilon}|^{2}\right]=\frac{1}{\varepsilon}\int_{0}^{1}\left(\mathbb{E}_{t}\left[e^{-\zeta(R_{T}+\lambda\xi_{T}^{t,\varepsilon}-R_{t})}|\xi_{T}^{t,\varepsilon}|^{2}\right](1-\lambda)\right)\mathrm{d}\lambda$
$\leq$ $\frac{1}{\varepsilon}e^{\zeta R_{t}}\left(\mathbb{E}_{t}\left[e^{-\zeta R_{T}}|\xi_{T}^{t,\varepsilon}|^{2}\right]\int_{0}^{1}(1-\lambda)^{2}\mathrm{d}\lambda+\mathbb{E}_{t}\left[e^{-\zeta(R_{T}+\xi_{T}^{t,\varepsilon})}|\xi_{T}^{t,\varepsilon}|^{2}\right]\int_{0}^{1}\lambda(1-\lambda)\mathrm{d}\lambda\right).$
Therefore, we only need to show that $\frac{1}{\varepsilon}\mathbb{E}_{t}\left[e^{-(R_{T}+\lambda\xi_{T}^{t,\varepsilon}-R_{t})}|\xi_{T}^{t,\varepsilon}|^{2}\right]\leq C|\eta|^{2}$ a.s. for $\lambda=0,1$, and for any $t\in[0,T)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$. By Lemma 2.7 and the Hölder’s inequality, there exists a constant $C$ only depending on $q,\|r\|_{\infty},\|\mu\|_{\infty},\|\sigma\|_{\infty},T,k,\|\pi\|_{\textup{BMO}}$ such that, for $\varepsilon$ sufficiently small,
$\frac{1}{\varepsilon}\mathbb{E}_{t}\left[e^{-\zeta(R_{T}+\lambda\xi_{T}^{t,\varepsilon}-R_{t})}|\xi_{T}^{t,\varepsilon}|^{2}\right]$ $\leq\frac{1}{\varepsilon}\left(\mathbb{E}_{t}\left[e^{-p\zeta(R_{T}-R_{t})}|\right]\right)^{\frac{1}{p}}\left(\mathbb{E}_{t}\left[|\xi_{T}^{t,\varepsilon}|^{2q}\right]\right)^{\frac{1}{q}}\left(\mathbb{E}_{t}\left[e^{-r\zeta\lambda\xi_{T}^{t,\varepsilon}}\right]\right)^{\frac{1}{r}}$
$\leq C\left(\mathbb{E}_{t}\left[e^{-p\zeta(R_{T}-R_{t})}|\right]\right)^{\frac{1}{p}}\left(\mathbb{E}_{t}\left[e^{-r\zeta\lambda\xi_{T}^{t,\varepsilon}}\right]\right)^{\frac{1}{r}}|\eta|^{2}$
$\leq C(\|\eta\|_{\infty})|\eta|^{2},\quad\lambda=0,1.$
Here, we choose $q,r$ such that $\frac{1}{p}+\frac{1}{q}+\frac{1}{r}=1$. Then, it holds that
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{0}^{1}e^{-\zeta(R_{T}+\lambda\xi_{T}^{t,\varepsilon}-R_{t})}(1-\lambda)d\lambda|\xi_{T}^{t,\varepsilon}|^{2}\right]\leq C(\|\eta\|_{\infty})|\eta|^{2}\quad a.s..$
Lemma 3.6.
For any admissible strategy $\pi\in\Pi_{0}$, consider the pair (Y,Y~)($\tilde{Y}$) and (Z,Z~)($\tilde{Z}$) in Lemma 3.4. We deduce the existence of a measurable set $E_{1}\subset[0,T)$ with $\operatorname{Leb}([0,T]\setminus E_{1})=0$ such that, for any $t\in E_{1}$, there exists a sequence $\{\varepsilon_{n}\}_{n\in\mathbb{N}}\subset(0,T-t)$ satisfying $\lim_{n\to\infty}\varepsilon_{n}^{t}=0$ and for any perturbation $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega;\mathbb{R})$, it holds that
$\lim\limits_{n\rightarrow\infty}\frac{1}{\varepsilon_{n}^{t}}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon_{n}^{t}}e^{-\zeta(R_{s}+Y_{s}-R_{t})}\left(a(s,\pi^{t,\varepsilon_{n}^{t}}_{s})-a(s,\pi_{s})-\zeta(Z_{s}^{1}+\sigma_{s}\pi_{s}+\rho Z_{s})\cdot\sigma_{s}\eta\right)\mathrm{d}s\right]$
$=$ $e^{-\zeta Y_{t}}\left((\mu_{t}-r_{t})\eta-\frac{1}{2}\sigma^{2}_{t}\eta^{2}-\sigma^{2}_{t}\eta\pi_{t}-\zeta(Z_{t}^{1}+\sigma_{t}\pi_{t}+\rho Z_{t})\cdot\sigma_{t}\eta\right),\quad a.s.$ (3.7)
and
$\lim\limits_{n\rightarrow\infty}\frac{1}{\varepsilon_{n}^{t}}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon_{n}^{t}}\left((\sigma_{s}\pi_{s}+\tilde{Z}^{1}_{s}+\rho\tilde{Z}_{s})^{2}-(\sigma_{s}\pi_{s}^{t,\varepsilon}+\tilde{Z}^{1}_{s}+\rho\tilde{Z}_{s})^{2}\right)\mathrm{d}s\right]$
$=$ $-2\sigma_{t}\eta(\sigma_{t}\pi_{t}+\tilde{Z}^{1}_{t}+\rho\tilde{Z}_{t}-\sigma_{t}^{2}\eta^{2}),\quad a.s..$ (3.8)
Proof.
Noting that $a(s,\pi^{t,\varepsilon_{n}^{t}}_{s})-a(s,\pi_{s})=(\mu_{s}-r_{s})\eta-\frac{1}{2}\sigma^{2}_{s}\eta^{2}-\sigma^{2}_{t}\eta\pi_{s}$ and $(\sigma_{s}\pi_{s}+\tilde{Z}^{1}_{s}+\rho\tilde{Z}_{s})^{2}-(\sigma_{s}\pi_{s}^{t,\varepsilon}+\tilde{Z}^{1}_{s}+\rho\tilde{Z}_{s})^{2}=-2\sigma_{s}\eta(\sigma_{s}\pi_{s}+\tilde{Z}^{1}_{s}+\rho\tilde{Z}_{s}-\sigma_{s}^{2}\eta^{2})$, we only need to prove that
$\lim\limits_{n\rightarrow\infty}\frac{1}{\varepsilon_{n}^{t}}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon_{n}^{t}}\alpha_{s}\left((\mu_{s}-r_{s})\eta-\frac{1}{2}\sigma^{2}_{s}\eta^{2}-\sigma^{2}_{t}\eta\pi_{s}-\zeta(Z_{s}^{1}+\sigma_{s}\pi_{s}+\rho Z_{s})\cdot\sigma_{s}\eta\right)\mathrm{d}s\right]$
$=$ $\alpha_{t}\left((\mu_{t}-r_{t})\eta-\frac{1}{2}\sigma^{2}_{t}\eta^{2}-\sigma^{2}_{t}\eta\pi_{t}-\zeta(Z_{t}^{1}+\sigma_{t}\pi_{t}+\rho Z_{t})\cdot\sigma_{t}\eta\right),\quad a.s.$
and
$\lim\limits_{n\rightarrow\infty}\frac{1}{\varepsilon_{n}^{t}}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon_{n}^{t}}2\sigma_{s}(\sigma_{s}\pi_{s}+\tilde{Z}^{1}_{s}+\rho\tilde{Z}_{s}-\sigma_{s}^{2}\eta^{2})\mathrm{d}s\right]=2\sigma_{t}(\sigma_{t}\pi_{t}+\tilde{Z}^{1}_{t}+\rho\tilde{Z}_{t}-\sigma_{t}^{2}\eta^{2}),\quad a.s..$
In view of $e^{-\zeta(R_{s}+Y_{s})}=\alpha_{s}\in L^{p}_{\mathbb{F}}(\Omega;C([0,T];\mathbb{R}))$ and the fact that $\mu-r$ and $\sigma$ are bounded, we have $\alpha(\mu-r),\alpha\sigma^{2}\in L^{p}(0,T;\mathbb{R})$ with $p>1$. It also holds that, for any $1<\gamma<p$ and $\frac{1}{\gamma}=\frac{1}{p}+\frac{1}{q^{\prime}}$,
$\mathbb{E}\left[\int_{0}^{T}|\alpha_{s}\pi_{s}|^{\gamma}\mathrm{d}s\right]\leq\left(\mathbb{E}\left[\sup\limits_{s\in[0,T]}|\alpha_{s}|^{p}\right]\right)^{\frac{\gamma}{p}}\left(\mathbb{E}\left[\left(\int_{0}^{T}|\pi_{s}|^{\gamma}\mathrm{d}s\right)^{\frac{q^{\prime}}{\gamma}}\right]\right)^{\frac{\gamma}{q^{\prime}}}<\infty,$
which implies that $\alpha\sigma^{2}\pi\in L^{\gamma}(0,T;\mathbb{R})$. In addition, $\alpha(Z^{1}+\sigma\pi)=\beta^{1},\;\alpha Z=\beta\in L^{p}_{\mathbb{F}}\left(\Omega;L^{2}(0,T;\mathbb{R})\right)\subset L^{p\wedge 2}(0,T;\mathbb{R})$ and $\tilde{\beta}^{1}=\tilde{Z}^{1}+\sigma\pi,\tilde{\beta}=\tilde{Z}\in L^{2}_{\mathbb{F}}\left(\Omega;L^{2}(0,T;\mathbb{R})\right)\subset L^{2}(0,T;\mathbb{R})$.
Therefore, by Lemma 3.3 in Hamaguchi_2021 , there exists a measurable set $E_{1}$ with $\operatorname{Leb}([0,T]\setminus E_{1})=0$ such that for any $t\in E_{1}$, we can choose a subsequence by recursively extracting subsequences and applying the diagonal selection argument that
$\lim\limits_{n\rightarrow\infty}\mathbb{E}_{t}\left[\frac{1}{\varepsilon_{n}^{t}}\int_{t}^{\varepsilon_{n}^{t}}P_{s}\right]=P_{t}.$
Here $P$ can be chosen to be $\alpha(\mu-r),\alpha\sigma^{2},\alpha\sigma^{2}\pi,\alpha(Z^{1}+\sigma\pi+\rho Z)$ and $\sigma(\sigma\pi+\tilde{Z}^{1}+\rho\tilde{Z}),\sigma^{3}\eta^{2}$. Then we obtain (3.6) and (3.6) with $E_{1}$ and the corresponding $\varepsilon_{n}^{t}$ for any $t\in E_{1}$, independent of $\eta$.
Remark 3.7.
Note that $a(s,\pi^{t,\varepsilon_{n}^{t}}_{s})-a(s,\pi_{s})$ depends on the perturbation strategy $\pi^{t,\varepsilon_{n}^{t}}$ only on $t,t+\varepsilon_{n}^{t}]$ and will be affected by $\eta$, we cannot directly apply Lemma 3.3 in Hamaguchi_2021. However, because the effect of $\eta$ can be separated, we can apply Lemma 3.3 to each element in two expressions and then use the diagonal selection argument to complete the proof.
Now we are ready to give the proof of Theorem 3.1.
Proof of Theorem 3.1.
Assume that $\hat{\pi}\in\Pi_{0}$ is an equilibrium strategy. Fix an arbitrary $\delta>0$ and take an arbitrary $t$ from the set $E_{1}$ in Lemma 3.6 . Let $\{\delta_{m}\}_{m\in\mathbb{N}}$ be a sequence such that $0<\delta_{m}\leq\delta,\;m\in\mathbb{N},$ and $\lim_{m\rightarrow\infty}\delta_{m}=0$. For each $m\in\mathbb{N}$, define $\eta_{m}$ by
$\eta_{m}=\begin{cases}\delta_{m}&e^{-\zeta Y_{t}}\theta-e^{-\zeta Y_{t}}\sigma_{t}\hat{\pi}_{t}-\zeta e^{-\zeta Y_{t}}(\sigma_{t}\hat{\pi}_{t}+Z_{t}^{1}+\rho Z_{t})-\gamma(\tilde{Z}^{1}_{t}+\sigma_{t}\hat{\pi}_{t}+\rho\tilde{Z}_{t})\geq 0,\\ -\delta_{m}&e^{-\zeta Y_{t}}\theta-e^{-\zeta Y_{t}}\sigma_{t}\hat{\pi}_{t}-\zeta e^{-\zeta Y_{t}}(\sigma_{t}\hat{\pi}_{t}+Z_{t}^{1}+\rho Z_{t})-\gamma(\tilde{Z}^{1}_{t}+\sigma_{t}\hat{\pi}_{t}+\rho\tilde{Z}_{t})<0.\end{cases}$
By the definition of the equilibrium, Lemma 3.6, Lemma 3.4 as well as Lemma 3.5, we obtain that there exists a sequence $\{\varepsilon_{n}^{t}\}_{n\in\mathbb{N}}\subset(0,T-t)$ satisfying $\lim_{n\to\infty}\varepsilon_{n}^{t}=0$ such that
[…]
$=$ $e^{-\zeta Y_{t}}\left((\mu_{t}-r_{t})\eta_{m}-\frac{1}{2}\sigma_{t}^{2}\eta_{m}^{2}-\sigma_{t}^{2}\hat{\pi}_{t}\eta_{m}-\zeta(Z_{t}^{1}+\sigma_{t}\hat{\pi}_{t}+\rho Z_{t})\cdot\sigma_{t}\eta_{m}\right)$
$\quad\quad-\frac{\gamma}{2}\sigma_{t}\eta_{m}\left(2\sigma_{t}\hat{\pi}+2\tilde{Z}^{1}_{t}+2\rho\tilde{Z}_{t}+\sigma_{t}\eta_{m}\right)-C(|\delta|)(|\eta_{m}|^{2}),\quad a.s..$
Dividing both sides of the above inequality by $\delta_{m}$, sending $m\rightarrow\infty$ and noting that the coefficient of $\sigma_{t}\eta_{m}$ tends to zero, we deduce that
$e^{-\zeta Y_{t}}\theta-e^{-\zeta Y_{t}}\sigma_{t}\hat{\pi}_{t}-\zeta e^{-\zeta Y_{t}}(\sigma_{t}\hat{\pi}_{t}+Z_{t}^{1}+\rho Z_{t})-\gamma(\tilde{Z}^{1}_{t}+\sigma_{t}\hat{\pi}_{t}+\rho\tilde{Z}_{t})=0.$
Consequently, the desired result holds that
$\sigma\hat{\pi}=\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}(Z^{1}+\rho Z)-\gamma(\tilde{Z}^{1}+\rho\tilde{Z})}{(\zeta+1)e^{-\zeta Y}+\gamma}.$
4 Study of BSDE System in the Incomplete Factor Model
In this section, we focus on the factor model and the asssociated BSDEs (3.1). The factor model provides an incomplete market setting with stochastic parameters in which the risk-free rate, the stock return rate, and the volatility rate can be expressed by a deterministic function of time $t$ and the factor $X_{t}$. It is assumed that the stochastic factor $X$ is governed by
$\mathrm{d}X_{t}=m(t,X_{t})\mathrm{d}t+\nu(t,X_{t})\mathrm{d}\bar{B_{t}},\quad X_{0}=x_{0}.$ (4.1)
It is assumed that $r_{t}=r(t,X_{t}),\mu_{t}=\mu(t,X_{t}),\sigma_{t}=\sigma(t,X_{t})$ for some functions $r,\mu,\sigma:[0,T]\times\mathbb{R}\to\mathbb{R}$ respectively. Moreover, we impose the following assumptions:
Assumption 4.1.
$\mu(t,x),r(t,x),\sigma(t,x)>0$ and $\theta(t,x)=\sigma^{-1}(\mu-r)(t,x)$ are bounded measurable (w.r.t. $\mathcal{B}([0,T]\times\mathbb{R})$) deterministic functions with respect to $(t,x)$. Here, $\theta$ represents the market price of risk, which is assumed to be positive and bounded.
Assumption 4.2.
The coefficients of the factor model satisfy that
1. The drift vector $m$ is uniformly bounded.
2. There exists a constant $\lambda>0$ such that $\lambda|z|^{2}\geq|z\nu(t,x)|^{2}\geq\frac{1}{\lambda}|z|^{2}$ for any $(t,x,z)\in[0,T]\times\mathbb{R}\times\mathbb{R}$.
3. There exists a constant $L$ such that
$|m(t,x)-m(t,x^{\prime})|+|\nu(t,x)-\nu(t,x^{\prime})|\leq L|x-x^{\prime}|.$
Under Assumption 4.2, SDE (4.1) admits a unique strong solution. It is noted that the log-return process $R$ does not appear in the BSDE system (3.1), which motivates us to consider the solution adapted to the filtration $\mathbb{F}_{\bar{B}}$. We can therefore conjecture that $Z^{1}=\tilde{Z}^{1}=0$ and aim to find the solution to the following Markovian BSDE system.
$\left\{\begin{aligned} &\mathrm{d}X_{s}=m(s,X_{s})\mathrm{d}t+\nu(s,X_{s})\mathrm{d}\bar{B}_{s},\\ &\mathrm{d}Y_{s}=\left(\frac{\zeta}{2}|\rho Z_{s}+\hat{u}_{s}|^{2}+\frac{\zeta(1-\rho^{2})}{2}|Z_{s}|^{2}-a(s,\hat{\pi}_{s})\right)\mathrm{d}s+Z_{s}\mathrm{d}\bar{B}_{s},\\ &\mathrm{d}\tilde{Y}_{s}=-a(s,\hat{\pi}_{s})\mathrm{d}s+\tilde{Z}_{s}\mathrm{d}\bar{B}_{s},\\ &Y_{T}=0,\quad\tilde{Y}_{T}=0,\\ &\hat{u}=\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}\rho Z-\gamma\rho\tilde{Z}}{(\zeta+1)e^{-\zeta Y}+\gamma},\hat{\pi}=\sigma^{-1}\hat{u},a(\cdot,\pi):=r+(\mu-r)\pi-\frac{1}{2}|\sigma\pi|^{2}.\end{aligned}\right$. (4.2)
4.1 Existence of time-consistent equilibrium when $\rho=0$
In this subsection, we consider the case $\rho=0$. Then it follows that
$\hat{\pi}_{t}=\sigma_{t}^{-1}\hat{u}_{t}=\sigma_{t}^{-1}\frac{e^{-\zeta Y_{t}}\theta(t,X_{t})}{(\zeta+1)e^{-\zeta Y_{t}}+\gamma},$ (4.3)
and the BSDE system (3.1) can be decoupled as
$\left\{\begin{aligned} &\mathrm{d}X_{s}=m(s,X_{s})\mathrm{d}t+\nu(s,X_{s})\mathrm{d}\bar{B}_{s},\\ &\mathrm{d}Y_{s}=\left(\frac{\zeta+1}{2}\hat{u}_{s}^{2}+\frac{\zeta}{2}|Z_{s}|^{2}-r(s,X_{s})-\theta(s,X_{s})\hat{u}_{s}\right)\mathrm{d}s+Z_{s}\mathrm{d}\bar{B}_{s},\\ &X_{0}=x,\quad Y_{T}=0.\end{aligned}\right$. (4.4)
and
$\left\{\begin{aligned} &\mathrm{d}\tilde{Y}_{s}=-r(s,X_{s})-\theta(s,X_{s})\hat{u}_{s}+\frac{1}{2}|\hat{u}_{t}|^{2}\mathrm{d}s+\tilde{Z}_{s}\mathrm{d}\bar{B}_{s},\\ &\tilde{Y}_{T}=0.\end{aligned}\right$. (4.5)
It is straightforward to see that one can first solve equation (4.4) and then substitute this solution into equation (4.5) to obtain the solution for the entire system. Regarding the existence of the solution to BSDE (4.4), we verify conditions (2A3), (3A2), (2A1), and (4A2) and apply Theorem 4.3 in FAN20161511, which leads to the following existence and uniqueness result.
Proposition 4.3.
If Assumptions 4.1 and 4.2 hold, the BSDE (4.4) has a unique solution $(Y,Z)\in\mathcal{S}^{\infty}\times L^{2}_{\mathbb{F}}(0,T;\mathbb{R})$.
By Proposition 4.3, the candidate equilibrium strategy defined in (4.3) satisfies $\hat{\pi}\in\mathcal{S}^{\infty}$. Therefore, BSDE (4.5) admits a unique solution $(\tilde{Y},\tilde{Z})\in\mathcal{S}^{\infty}\times L^{2}_{\mathbb{F}}(0,T;\mathbb{R})$ and $\tilde{Y}$ is given by
$\tilde{Y}_{t}=\mathbb{E}_{t}\left[\int_{t}^{T}\left(r(s,X_{s})+\theta(s,X_{s})\hat{u}_{s}-\frac{1}{2}|\hat{u}_{s}|^{2}\right)\mathrm{d}s\right].$
In the following, we establish that $\hat{\pi}$ is an equilibrium strategy by means of BSDEs (4.5) and (4.3), along with Lemma 3.4. This result is stated in the theorem below.
Theorem 4.4.
When $\rho=0$, suppose that Assumptions 4.1 and 4.2 hold, and let $Y$ denote the bounded solution to BSDE (4.4). Then the strategy defined in (4.3) is an equilibrium strategy.
Proof.
For notational simplicity, let $\hat{R}=R^{\hat{\pi}}$ denote the log-return process associated with $\hat{\pi}$. Moreover, we use $\pi^{t,\varepsilon}$ and $\xi^{t,\varepsilon}$ to represent the perturbation strategy $\hat{\pi}^{t,\varepsilon,\eta}$ and the corresponding difference process for any fixed $t\in[0,T)$, $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R}^{d})$ and $\varepsilon\in(0,T-t)$.
We first verify that $\hat{\pi}$ is an admissible strategy. Since $\hat{\pi}$ is bounded, it follows that $\hat{\pi}\in H_{\textup{BMO}}$. It is straightforward to show that
$\mathbb{E}\left[\exp\left(c\left|\int_{t}^{T}|\sigma_{s}\hat{\pi}_{s}|\mathrm{d}B_{s}\right|\right)\right]\leq 2\exp\left(\frac{c^{2}}{2}\|\sigma\hat{\pi}\|_{\infty}^{2}(T-t)\right),\quad\forall c>0.$
Combining this with $\hat{R}_{T}-\hat{R}_{t}=\int_{t}^{T}\left(r_{s}+(\mu_{s}-r_{s})\hat{\pi}_{s}-\frac{1}{2}|\sigma_{s}\hat{\pi}_{s}|^{2}\right)\mathrm{d}s+\sigma_{s}\hat{\pi}_{s}\mathrm{d}B_{s}$, we conclude that $\mathbb{E}_{t}\left[\exp\left(-p\zeta(R^{\hat{\pi}}_{T}-R^{\hat{\pi}}_{t})\right)\right]<\infty,\,a.s.$ for any $p>1$. Thus, $\hat{\pi}$ is an admissible strategy.
Next, we show that $\hat{\pi}$ is indeed an equilibrium strategy. By Lemma 3.4 and $Z^{1}=\tilde{Z}^{1}=0$, it holds that
[…]
$+\frac{\gamma}{2}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}((\sigma_{s}\hat{\pi}_{s})^{2}-(\sigma_{s}\pi_{s}^{t,\varepsilon})^{2})\mathrm{d}s\right]$
$=$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta(\hat{R}_{s}+Y_{s}-\hat{R}_{t})}\left((\mu_{s}-r_{s})\eta-\frac{1}{2}\sigma_{s}^{2}\eta^{2}-(\zeta+1)\sigma_{s}^{2}\hat{\pi}_{s}\eta\right)\mathrm{d}s\right]$
$-\frac{\gamma}{2}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\sigma_{s}\eta\left(2\sigma_{s}\hat{\pi}_{s}+\sigma_{s}\eta\right)\mathrm{d}s\right].$
We first prove that
$\lim\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left|\left(e^{-\zeta\left(Y_{s}+\hat{R}_{s}-\hat{R}_{t}\right)}-e^{-\zeta Y_{s}}\right)\left((\mu_{s}-r_{s})\eta-\frac{1}{2}\sigma_{s}^{2}\eta^{2}-(\zeta+1)\sigma_{s}^{2}\hat{\pi}_{s}\eta\right)\right|\mathrm{d}s\right]=0,\quad a.s..$
Since $r$, $\mu$, $\sigma$, $\hat{\pi}$, and $Y$ are all bounded, it suffices to show that
$\lim\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left|\left(e^{-\zeta\left(\hat{R}_{s}-\hat{R}_{t}\right)}-1\right)\right|\mathrm{d}s\right]=0,\quad a.s.,$
which follows immediately from the fact that $\{e^{-\zeta\left(\hat{R}_{s}-\hat{R}_{t}\right)}\}_{s\in[t,T]}\in L_{\mathbb{F}}^{p}\left(\Omega;\left(C[t,T];\mathbb{R}\right)\right)$. Therefore
[…]
$\leq$ $\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\bigg(e^{-\zeta Y_{s}}\Big((\mu_{s}-r_{s})\eta-\frac{1}{2}\sigma_{s}^{2}\eta^{2}-(\zeta+1)\sigma_{s}\eta\hat{u}_{s})\Big)-\frac{\gamma}{2}\sigma_{s}^{2}\eta^{2}-\gamma\sigma_{s}\eta\hat{u}_{s}\bigg)\mathrm{d}s\right]$
$=$ $\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left(-\frac{1}{2}\left(e^{-\zeta Y_{s}}+\gamma\right)(\sigma_{s}\eta)^{2}+\sigma_{s}\eta\left[e^{-\zeta Y_{s}}\left(\theta_{s}-(\zeta+1)\hat{u}_{s}\right)-\gamma\hat{u}_{s}\right]\right)\mathrm{d}s\right].$
Substituting $\hat{u}=\frac{e^{-\zeta Y}\theta}{(\zeta+1)e^{-\zeta Y}+\gamma}$ into the above expression yields that
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\left(J(t,\hat{R}_{t},\pi^{t,\varepsilon})-J(t,\hat{R}_{t},\hat{\pi})\right)\leq\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left(-\frac{1}{2}\left(e^{-\zeta Y_{s}}+\gamma\right)(\sigma_{s}\eta)^{2}\right)\mathrm{d}s\right]\leq 0.$
Therefore, $\hat{\pi}$ is a time-consistent equilibrium strategy. ∎
4.2 Existence of time-consistent equilibrium under trading constraint
In this subsection, we assume that the trading strategy $\pi_{t}$ takes values in a bounded convex Borel set $A$. Consequently, $u_{t}=\sigma_{t}\pi_{t}$ is also bounded. To be precise, we introduce the following definition.
Definition 4.5 ((Admissible trading strategy with constraint $A$)).
A trading strategy $\pi$ is said to be admissible with constraint $A$ if $\pi_{t}\in A$ a.s. for any $t\in[0,T]$. We denote by $\Pi_{A}$ the set of all admissible trading strategies with constraint $A$.
As in the proof of Theorem 4.4, it is straightforward that $\Pi_{A}\subseteq\Pi_{0}$. Following yan2019time and liang2025integral, we next introduce the modified definitions of the perturbation strategy and the equilibrium strategy.
Definition 4.6.
A strategy $\hat{\pi}\in\Pi_{A}$ is called an equilibrium strategy if
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{J(t,\hat{\pi}^{t,\varepsilon,a})-J(t,\hat{\pi})}{\varepsilon}\leq 0\quad a.s.,$
for any $t\in[0,T)$ and any $a\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,A)$ . Here
$\hat{\pi}^{t,\varepsilon,a}=\left\{\begin{aligned} &a,\quad t\in[t,t+\varepsilon),\\ &\hat{\pi},\quad t\in[t+\varepsilon,T],\end{aligned}\right$.
denotes the modified perturbation strategy.
Remark 4.7.
In this modified definition, instead of adding an $\mathcal{F}_{t}$-measurable random variable $\eta$ to the original strategy $\hat{\pi}$, we perturb the strategy by directly replacing it with an $\mathcal{F}_{t}$-measurable random variable $a$ taking values in $A$. This modification is made because requiring the perturbed strategy to remain admissible, i.e., $\pi_{s}+\eta\in A$ for $s\in[t,t+\varepsilon)$, would impose a strong restriction on $\eta$, possibly making it impossible to perturb many strategies.
We first introduce a lemma for the projection onto a convex closed set.
Lemma 4.8.
Let $U\subset\mathbb{R}^{n}$ be a convex and closed set. Fix $w\in\mathbb{R}^{n}$, then we have the following two assertions: (i) if $u=P_{U}(w)$ is the orthogonal projection of $w$ onto $U$. Let $h\in\mathbb{R}^{n}$ satisfy $u+h\in U$. Then for any $\alpha>0$
$\bigl|\alpha(w-u)-h\bigr|\geq\alpha|w-u|,$ (4.6)
and equality can occur only in the trivial case $h=0$; (ii) if $U\ni u\neq P_{U}(w)$, then for any $\alpha>0$, there exists an $h$ such that $u+h\in U$ and
$\bigl|\alpha(w-u)-h\bigr|<\alpha|w-u|.$
Moreover, $h$ can be chosen to be $\lambda\left(P_{U}(w)-u\right)$ for any $\lambda\in(0,2\alpha)$.
Proof.
We first prove assertion (i). By the characterization of the orthogonal projection onto a convex set, for every $z\in U$, we have
$\langle w-u,\,z-u\rangle\leq 0.$
Choosing $z=u+h\in U$ yields
$\langle w-u,\,h\rangle\leq 0.$ (4.7)
It then follows that
$\bigl|\alpha(w-u)-h\bigr|^{2}=\alpha^{2}|w-u|^{2}+|h|^{2}-2\alpha\langle w-u,\,h\rangle\geq\alpha^{2}|w-u|^{2}.$
Taking the square root yields the desired inequality (4.6). Moreover, the equality requires both $\langle w-u,h\rangle=0$ and $|h|^{2}=0$, which implies that $h=0$.
Next we verify assertion (ii). As $u$ is not the projection, the vector $v=P_{U}(w)$ satisfies that
$\langle w-u,\,v-u\rangle\geq|v-u|^{2}>0.$
Let $h=\lambda(v-u),\;\lambda\in(0,1]$ ,then
$\bigl|\alpha(w-u)-h\bigr|<\alpha|w-u|\iff 2\alpha\langle w-u,h\rangle>|h|^{2}\Longleftarrow 0<\lambda<2\alpha.$
Since $\alpha>0$, the above inequality holds for sufficiently small $\lambda>0$. By the convexity of $U$, $u+h\in U$, which completes the proof. ∎
Remark 4.9.
In this lemma, we consider a convex and closed set in $\mathbb{R}^{n}$ for any positive integer $n$, rather than in $\mathbb{R}$, thereby showing that our subsequent proof remains valid in the presence of multiple risky assets.
By an analogous argument in section 3, we have the following necessary condition for an equilibrium strategy.
Theorem 4.10 (Necessary condition).
[…]
$\hat{\pi}=\sigma^{-1}\mathcal{P}_{\sigma A}\left(\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}(Z^{1}+\rho Z)-\gamma(\tilde{Z}^{1}+\rho\tilde{Z})}{(\zeta+1)e^{-\zeta Y}+\gamma}\right),\quad t\in[0,T).$ (4.8)
Proof.
[…]
$=$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta(\hat{R}_{s}+Y_{s}-\hat{R}_{t})}(a(s,a)-a(s,\hat{\pi}_{s})-\zeta\sigma_{s}\hat{\pi}_{s}\cdot\Delta_{s}-\zeta\Delta_{s}\left(\tilde{Z}^{1}_{s}+\rho Z_{s}\right))\mathrm{d}s\right]$
$-\zeta\mathbb{E}_{t}\left[\int_{0}^{1}e^{-\zeta(\hat{R}_{T}+\lambda\xi_{T}^{t,\varepsilon,a}-\hat{R}_{t})}(1-\lambda)d\lambda|\xi_{T}^{t,\varepsilon,a}|^{2}\right]$
$+\frac{\gamma}{2}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left((\sigma_{s}\hat{\pi}_{s}+\tilde{Z}^{1}_{s}+\rho\tilde{Z}_{s})^{2}-(\sigma_{s}a+\tilde{Z}^{1}_{s}+\rho\tilde{Z}_{s})^{2}\right)\mathrm{d}s\right].$
Moreover, by the same argument in Lemma 3.5 and Lemma 3.6, there exists a sequence $\{\varepsilon_{n}^{t}\}$ and a fixed constant $C$ independent of $a$ such that for $a.e.$ $t\in[0,T)$,
$\lim\limits_{n\rightarrow\infty}\frac{1}{\varepsilon_{n}^{t}}\mathbb{E}_{t}$ $\left[\int_{t}^{t+\varepsilon_{n}^{t}}e^{-\zeta(\hat{R}_{s}+Y_{s}-\hat{R}_{t})}\left(a(s,a)-a(s,\hat{\pi}_{s})-\zeta(Z_{s}^{1}+\sigma_{s}\hat{\pi}_{s}+\rho Z_{s})\cdot\Delta_{s}\right)\mathrm{d}s\right]$
$=$ $e^{-\zeta Y_{t}}\left(\theta_{t}\Delta_{t}-\frac{1}{2}\Delta_{t}^{2}-\sigma_{t}\pi_{t}\Delta_{t}-\zeta(Z_{t}^{1}+\sigma_{t}\hat{\pi}_{t}+\rho Z_{t})\cdot\Delta_{t}\right),\quad a.s.,$
$\lim\limits_{n\rightarrow\infty}\frac{1}{\varepsilon_{n}^{t}}\mathbb{E}_{t}\left[\int_{0}^{1}e^{-\zeta(\hat{R}_{T}+\lambda\xi_{T}^{t,\varepsilon,a}-\hat{R}_{t})}(1-\lambda)d\lambda|\xi_{T}^{t,\varepsilon,a}|^{2}\right]\leq C\Delta_{t}^{2},\quad a.s.,$
and
$\lim\limits_{n\rightarrow\infty}\frac{1}{\varepsilon_{n}^{t}}\mathbb{E}_{t}$ $\left[\int_{t}^{t+\varepsilon_{n}^{t}}\left((\sigma_{s}\hat{\pi}_{s}+\tilde{Z}^{1}_{s}+\rho\tilde{Z}_{s})^{2}-(\sigma_{s}a+\tilde{Z}^{1}_{s}+\rho\tilde{Z}_{s})^{2}\right)\mathrm{d}s\right]$
$=$ $(\sigma_{t}\hat{\pi}_{t}+\tilde{Z}^{1}_{t}+\rho\tilde{Z}_{t})^{2}-(\sigma_{t}a+\tilde{Z}^{1}_{t}+\rho\tilde{Z}_{t})^{2},\quad a.s..$
Then the equilibrium strategy $\hat{\pi}$ should satisfy that
$-\frac{e^{-\zeta Y_{t}}+\gamma}{2}(\sigma_{t}a)^{2}+\left(e^{-\zeta Y_{t}}\left(\theta_{t}-\zeta\left(\sigma_{t}\hat{\pi}_{t}+Z^{1}_{t}+\rho Z_{t}\right)\right)-\gamma\left(\tilde{Z}^{1}_{t}+\rho\tilde{Z}_{t}\right)\right)\sigma_{t}a$
$\leq$ $-\frac{e^{-\zeta Y_{t}}+\gamma}{2}(\sigma_{t}\hat{\pi}_{t})^{2}+\left(e^{-\zeta Y_{t}}\left(\theta_{t}-\zeta\left(\sigma_{t}\hat{\pi}_{t}+Z^{1}_{t}+\rho Z_{t}\right)\right)-\gamma\left(\tilde{Z}^{1}_{t}+\rho\tilde{Z}_{t}\right)\right)\sigma_{t}\hat{\pi}_{t}+C\Delta_{t}^{2},\quad a.s..$ (4.10)
It follows that
$\frac{e^{-\zeta Y_{t}}+\gamma}{2}\Delta_{t}^{2}+\left(e^{-\zeta Y_{t}}\left(\theta_{t}-(\zeta+1)\sigma_{t}\hat{\pi}_{t}-\zeta\left(Z_{t}^{1}+\rho Z_{t}\right)\right)-\gamma\left(\tilde{Z}_{t}^{1}+\rho\tilde{Z}\right)-\gamma\sigma_{t}\hat{\pi}_{t}\right)\Delta_{t}\leq 0,\quad a.s..$ (4.11)
Suppose, to the contrary, that there exists some $a\in L_{\mathcal{F}_{t}}(\Omega,A)$ such that (4.11) fails to hold. Define a new admissible strategy $a_{\lambda}=\hat{\pi}_{t}+\lambda\sigma_{t}^{-1}\Delta\in A$ with sufficiently small constant $\lambda\in(0,\frac{\gamma}{\gamma+2C})$. On one hand, by (4.2), we have
$0\geq$ $-\frac{e^{-\zeta Y_{t}}+\gamma}{2}(\sigma_{t}a_{\lambda})^{2}+\left(e^{-\zeta Y_{t}}\left(\theta_{t}-\zeta\left(\sigma_{t}\hat{\pi}_{t}+Z^{1}_{t}+\rho Z_{t}\right)\right)-\gamma\left(\tilde{Z}^{1}_{t}+\rho\tilde{Z}_{t}\right)\right)\sigma_{t}a_{\lambda}-C|a_{\lambda}-\hat{\pi}_{t}|^{2}$
$+\frac{e^{-\zeta Y_{t}}+\gamma}{2}(\sigma_{t}\hat{\pi}_{t})^{2}-\left(e^{-\zeta Y_{t}}\left(\theta_{t}-\zeta\left(\sigma_{t}\hat{\pi}_{t}+Z^{1}_{t}+\rho Z_{t}\right)\right)-\gamma\left(\tilde{Z}^{1}_{t}+\rho\tilde{Z}_{t}\right)\right)\sigma_{t}\hat{\pi}_{t}$
$=$ $\lambda\left(-\frac{e^{-\zeta Y_{t}}+\gamma}{2}\lambda\Delta_{t}^{2}+\left(e^{-\zeta Y_{t}}\left(\theta_{t}-(\zeta+1)\sigma_{t}\hat{\pi}_{t}-\zeta\rho Z_{t}\right)-\gamma\rho\tilde{Z}-\gamma\sigma_{t}\hat{\pi}_{t}\right)\Delta_{t}-C\lambda\Delta_{t}^{2}\right),\quad a.s..$
On the other hand, as $\lambda\in(0,\frac{\gamma}{\gamma+2C})$ and (4.11) does not hold for $a$, we have
$\lambda\left(-\frac{e^{-\zeta Y_{t}}+\gamma}{2}\lambda\Delta_{t}^{2}+\left(e^{-\zeta Y_{t}}\left(\theta_{t}-(\zeta+1)\sigma_{t}\hat{\pi}_{t}-\zeta\rho Z_{t}\right)-\gamma\rho\tilde{Z}-\gamma\sigma_{t}\hat{\pi}_{t}\right)\Delta_{t}-C\lambda\Delta_{t}^{2}\right)$
$\geq$ $\lambda\left(-\frac{e^{-\zeta Y_{t}}+\gamma}{2}\Delta_{t}^{2}+\left(e^{-\zeta Y_{t}}\left(\theta_{t}-(\zeta+1)\sigma_{t}\hat{\pi}_{t}-\zeta\rho Z_{t}\right)-\gamma\rho\tilde{Z}-\gamma\sigma_{t}\hat{\pi}_{t}\right)\Delta_{t}\right)>0,$
on a set of positive measure, which leads to a contradiction. Therefore, (4.11) must hold for all $a\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,A)$. Consequently, we obtain
$\left|\Delta_{t}-\frac{e^{-\zeta Y_{t}}\left(\theta_{t}-(\zeta+1)\sigma_{t}\hat{\pi}_{t}-\zeta\left(Z_{t}^{1}+\rho Z_{t}\right)\right)-\gamma\left(\tilde{Z}_{t}^{1}+\rho\tilde{Z}\right)-\gamma\sigma_{t}\hat{\pi}_{t}}{e^{-\zeta Y_{t}}+\gamma}\right|$
$\geq$ $\left|\frac{e^{-\zeta Y_{t}}\left(\theta_{t}-(\zeta+1)\sigma_{t}\hat{\pi}_{t}-\zeta\left(Z_{t}^{1}+\rho Z_{t}\right)\right)-\gamma\left(\tilde{Z}_{t}^{1}+\rho\tilde{Z}\right)-\gamma\sigma_{t}\hat{\pi}_{t}}{e^{-\zeta Y_{t}}+\gamma}\right|,\quad a.s..$
Take $\alpha=\frac{(\zeta+1)e^{-\zeta Y_{t}}+\gamma}{e^{-\zeta Y_{t}}+\gamma}$ and $w=\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}(Z^{1}+\rho Z)-\gamma(\tilde{Z}^{1}+\rho\tilde{Z})}{(\zeta+1)e^{-\zeta Y}+\gamma}$, the above inequality can be converted to
$\left|\sigma_{t}a-\hat{u}_{t}-\alpha_{t}\left(w_{t}-\hat{u}_{t}\right)\right|\geq\left|\alpha_{t}\left(w_{t}-\hat{u}_{t}\right)\right|,\quad a.s..$
By Lemma 4.8(ii), if $\hat{u}_{t}\neq\mathcal{P}_{\sigma_{t}A}\left(w_{t}\right)$, as $\alpha_{t}>1$, we can take $\lambda=1$ and then
$\left|\mathcal{P}_{\sigma_{t}A}\left(w_{t}\right)-\hat{u}_{t}-\alpha_{t}(w_{t}-\hat{u}_{t})\right|<\left|\alpha_{t}(w_{t}-\hat{u}_{t})\right|.$
Thus we know that $\hat{u}\neq\mathcal{P}_{\sigma_{t}A}\left(w_{t}\right)$ can not be an equilibrium strategy by letting $\sigma_{t}a=\mathcal{P}_{\sigma_{t}A}\left(w_{t}\right)$. Therefore, (4.8) holds. ∎
As $R$ does not appear in the BSDE system, we also assume that $Z^{1}=\tilde{Z}^{1}=0$
$\left\{\begin{aligned} &\mathrm{d}X_{s}=m(s,X_{s})\mathrm{d}t+v(s,X_{s})\mathrm{d}\bar{B}_{s},\\ &\mathrm{d}Y_{s}=\left(\frac{\zeta}{2}|\rho Z_{s}+\hat{u}_{s}|^{2}+\frac{\zeta(1-\rho^{2})}{2}|Z_{s}|^{2}-a(s,\hat{\pi}_{s})\right)\mathrm{d}s+Z_{s}\mathrm{d}\bar{B}_{s},\\ &\mathrm{d}\tilde{Y}_{s}=-a(s,\hat{\pi}_{s})\mathrm{d}s+\tilde{Z}_{s}\mathrm{d}\bar{B}_{s},\\ &Y_{T}=0,\quad\tilde{Y}_{T}=0,\quad\hat{u}=\sigma\hat{\pi}=\mathcal{P}_{\sigma_{t}A}\left(\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}\rho Z-\gamma\rho\tilde{Z}}{(\zeta+1)e^{-\zeta Y}+\gamma}\right).\end{aligned}\right$. (4.12)
To establish the existence of the BSDE system, we only need to verify the AB condition and the BF condition (see their definitions in Appendix B).
AB condition:
The generator of $Y$ satisfy that
[…]
The generator of $\tilde{Y}$ satisfy that
$r-\theta^{2}\leq-r-\theta\hat{u}+\frac{1}{2}\hat{u}^{2}\leq\frac{1}{2}(\hat{u}-\theta)^{2}$
Let $a_{i},i=1,2,3,4$ be (1,0),(1,0),(0,1)(0,1)(1,0),(-1,0),(0,1)(0,-1). Then $a_{i},i=1,2,3,4$ can positively span $\mathbb{R}^{2}$ and the generator satisfy the AB condition.
BF condition
The generator of $Y$ satisfy that
$\left|\frac{\zeta}{2}Z^{2}+\zeta\rho Z\hat{u}-r-\theta\hat{u}+\frac{\zeta+1}{2}\hat{u}^{2}\right|\leq\frac{\zeta(1+\rho^{2})}{2}Z^{2}+k(t),$
with $k(t)=r(t,X_{t})+\theta(t,X_{t})\sup\limits_{\hat{u}\in\sigma_{t}A}|\hat{u}|+\frac{2\zeta+1}{2}\sup\limits_{\hat{u}\in\sigma_{t}A}|\hat{u}|^{2}$.
Let $C_{n}=\sup\limits_{t\in[0,T]}\left\{|\theta(t,X_{t})|+k(t)\right\}$ and the generator of $\tilde{Y}$ satisfy that
$|-r-\theta\hat{u}+\frac{1}{2}\hat{u}^{2}|\leq k(t)$
Then the BSDE system under the bounded equilibrium strategy admits a locally Holder and BMO solution (see Theorem 2.14 in xing_class_2018). Moreover, $Y$ and $\tilde{Y}$ are bounded due to the terminal condition $Y_{T}=0$.
Theorem 4.11.
Under the trading constraint $u_{t}=\sigma_{t}\pi_{t}\in\sigma_{t}A$, where $A\subset\mathbb{R}$ is a bounded closed convex Borel set, assume that (Y,Y~,Z,Z~)($\tilde{Y},$$\tilde{Z}$) is a bounded BMO solution to the BSDE (4.12), then
$\hat{u}_{t}=\sigma_{t}\hat{\pi}_{t}=\mathcal{P}_{\sigma_{t}A}\left(\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}\rho Z-\gamma\rho\tilde{Z}}{(\zeta+1)e^{-\zeta Y}+\gamma}\right)$ (4.13)
is an equilibrium strategy.
Proof.
It is clear that $\hat{\pi}\in\Pi_{A}$. We now prove that $\hat{\pi}$ is an equilibrium strategy. Fix $t\in[0,T)$ and let $a\in L_{\mathcal{F}_{t}}^{\infty}(\Omega;A)$; denote the corresponding perturbed strategy by $\pi^{t,\varepsilon}$. Put $\Delta_{s}=\sigma_{s}(a-\hat{\pi}_{s})$. Then, by (4.9), it holds almost surely that
[…]
$\leq$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta(\hat{R}_{s}+Y_{s}-\hat{R}_{t})}(a(s,a)-a(s,\hat{\pi}_{s})-\zeta\sigma_{s}\hat{\pi}_{s}\cdot\Delta_{s}-\zeta\rho\Delta_{s}Z_{s})\mathrm{d}s\right]$
$+\frac{\gamma}{2}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}((\sigma_{s}\hat{\pi}_{s}+\rho\tilde{Z}_{s})^{2}-(\sigma_{s}a+\rho\tilde{Z}_{s})^{2})\mathrm{d}s\right]$
$=$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta(\hat{R}_{s}+Y_{s}-\hat{R}_{t})}\left(\theta_{s}\Delta_{s}-\frac{1}{2}\Delta_{s}^{2}-(\zeta+1)\sigma_{s}\hat{\pi}_{s}\Delta_{s}-\zeta\rho\Delta_{s}Z_{s}\right)\mathrm{d}s\right]$
$-\frac{\gamma}{2}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\Delta_{s}\left(2\sigma_{s}\hat{\pi}_{s}+2\rho\tilde{Z}_{s}+\Delta_{s}\right)\mathrm{d}s\right].$
To verify that the candidate strategy is a time-consistent equilibrium, we need to show that
$\lim\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left|\left(e^{-\zeta\left(Y_{s}+\hat{R}_{s}-\hat{R}_{t}\right)}-e^{-\zeta Y_{s}}\right)\left(\theta_{s}\Delta_{s}-\frac{1}{2}\Delta_{s}^{2}-(\zeta+1)\sigma_{s}\hat{\pi}_{s}\Delta_{s}-\zeta\rho\Delta_{s}Z_{s}\right)\right|\mathrm{d}s\right]=0.$
It suffices to prove that
$\lim\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left|\left(e^{-\zeta\left(\hat{R}_{s}-\hat{R}_{t}\right)}-1\right)Z_{s}\right|\mathrm{d}s\right]=0,\quad a.s.,$
Note that $\left\{e^{-\zeta(\hat{R}_{s}-\hat{R}_{t})}\right\}_{s\in[t,T]}$ satisfies
$\mathrm{d}\left(e^{-\zeta(\hat{R}_{s}-\hat{R}_{t})}\right)=-\zeta e^{-\zeta(\hat{R}_{s}-\hat{R}_{t})}\left(\left(a(s,\hat{\pi}_{s})-\frac{\zeta}{2}\sigma_{s}^{2}\hat{\pi}_{s}^{2}\right)\mathrm{d}s+\sigma_{s}\hat{\pi}_{s}\mathrm{d}B_{s}\right).$
Because $\hat{\pi}$ is bounded, by Theorem 3.4.3 in zhang_backward_2017 and the same argument of Lemma 2.7, we deduce that
$\mathbb{E}_{t}\left[\sup_{s\in[t,t+\varepsilon]}\left|e^{-\zeta(\hat{R}_{s}-\hat{R}_{t})}-1\right|^{2}\right]\leq C\mathbb{E}\left[\left(\int_{t}^{t+\varepsilon}\left|a(s,\hat{\pi}_{s})-\frac{\zeta}{2}\sigma_{s}^{2}\hat{\pi}_{s}^{2}\right|\mathrm{d}s\right)^{2}+\int_{t}^{t+\varepsilon}\sigma_{s}^{2}\hat{\pi}_{s}^{2}\mathrm{d}s\right]\leq C\varepsilon.$
Moreover, since $Z\in H_{\textup{BMO}}$, $\int_{t}^{T}|Z_{s}|^{2}\mathrm{d}s<\infty,\quad a.s.$ and for fixed $t$ $\int_{t}^{t+\varepsilon}|Z_{s}|^{2}\mathrm{d}s\;\rightarrow\;0\quad a.s..$ By the Conditional Dominated Convergence Theorem
$\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}|Z_{s}|^{2}\mathrm{d}s\right]\;\longrightarrow\;0\quad a.s..$
Then
$\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left|\left(e^{-\zeta\left(\hat{R}_{s}-\hat{R}_{t}\right)}-1\right)Z_{s}\mathrm{d}s\right|\right]$
$\leq$ $\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left|e^{-\zeta\left(\hat{R}_{s}-\hat{R}_{t}\right)}-1\right|^{2}\mathrm{d}s\right]^{\frac{1}{2}}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}Z_{s}^{2}\mathrm{d}s\right]^{\frac{1}{2}}$
$\leq$ $\frac{1}{\sqrt{\varepsilon}}\mathbb{E}_{t}\left[\sup_{s\in[t,t+\varepsilon]}\left|e^{-\zeta(\hat{R}_{s}-\hat{R}_{t})}-1\right|^{2}\right]^{\frac{1}{2}}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}Z_{s}^{2}\mathrm{d}s\right]^{\frac{1}{2}}$
$\leq$ $C\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}Z_{s}^{2}\mathrm{d}s\right]^{\frac{1}{2}}\;\longrightarrow 0,\quad a.s..$
Similar to Theorem 4.4, one can derive that
[…]
$\leq$ $\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta Y_{s}}\left(\theta_{s}\Delta_{s}-\frac{1}{2}\Delta_{s}^{2}-(\zeta+1)\sigma_{s}\eta\hat{u}_{s}-\zeta\rho\Delta_{s}Z_{s}\right)\right$.
$\left.-\frac{\gamma}{2}\Delta_{s}^{2}-\gamma\Delta_{s}(\hat{u}_{s}+\rho\tilde{Z}_{s})\mathrm{d}s\right]$
[…]
$\left.\left.+\Delta_{s}\left[e^{-\zeta Y_{s}}\left(\theta_{s}-(\zeta+1)\hat{u}_{s}-\zeta\rho Z_{s}\right)-\gamma(\hat{u}_{s}+\rho\tilde{Z}_{s})\right]\right)\mathrm{d}s\right].$
As $\sigma>0$ and $\sigma_{s}A$ is a closed and convex set, by Lemma 4.8(i), for any $\Delta_{s}$ such that $\hat{u}_{s}+\Delta_{s}\in\sigma_{s}A$, it holds that
$\left|\frac{(\zeta+1)e^{-Y_{s}}+\gamma}{e^{-Y_{s}}+\gamma}\left(\frac{e^{-\zeta Y_{s}}\theta_{s}-\zeta e^{-\zeta Y_{s}}\rho Z_{s}-\gamma\rho\tilde{Z}_{s}}{(\zeta+1)e^{-Y_{s}}+\gamma}-\hat{u}_{s}\right)-\Delta_{s}\right|$
$\geq$ $\left|\frac{(\zeta+1)e^{-Y_{s}}+\gamma}{e^{-Y_{s}}+\gamma}\left(\frac{e^{-\zeta Y_{s}}\theta_{s}-\zeta e^{-\zeta Y_{s}}\rho Z_{s}-\gamma\rho\tilde{Z}_{s}}{(\zeta+1)e^{-Y_{s}}+\gamma}-\hat{u}_{s}\right)\right|.$
Thus $\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\left(J(t,\hat{R}_{t},\pi^{t,\varepsilon,a})-J(t,\hat{R}_{t},\hat{\pi})\right)\leq 0$ , a.s., which completes the proof. ∎
5 Approximate Time-Consistent Equilibrium for Small $\rho\neq 0$
The goal of this section is to construct and verify an approximate time-consistent equilibrium in the general case with small $\rho\neq 0$, where the approximation error can be shown to be the order $O(\rho^{2})$. Let us first introduce the definition of an approximate time-consistent equilibrium strategy.
Definition 5.1.
For $\hat{\pi}\in\Pi_{0}$, $\hat{\pi}$ is called an approximate time-consistent equilibrium strategy with an approximation error $M$ if
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{J(t,\hat{\pi}^{t,\varepsilon,\eta})-J(t,\hat{\pi})}{\varepsilon}\leq M\quad a.s.,$
for any $t\in[0,T)$ and any $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$. Moreover, if $M=O(|\rho|^{\alpha})$, we say that the approximation error [^5] The order of the approximation error effectively characterizes the convergence rate of the equilibrium approximation as $\rho$ tends to 0. is of the order $O(|\rho|^{\alpha})$.
Recall that for $\rho=0$, the equilibrium strategy follows the form of $\hat{u}=\frac{e^{-\zeta Y}\theta}{(\zeta+1)e^{-\zeta Y}+\gamma}$, where $Y$ is the solution to the BSDE (4.4). For the case $\rho\neq 0$, we consider a trading strategy of the same structure, except that the process $Y$ is replaced by the solution $Y^{\rho}$ to the following BSDE, Specifically, let $Y^{\rho}$ and $\tilde{Y}^{\rho}$ satisfy
$\left\{\begin{aligned} &\mathrm{d}X_{s}=m(s,X_{s})\mathrm{d}t+v(s,X_{s})\mathrm{d}\bar{B}_{s},\\ &\mathrm{d}Y^{\rho}_{s}=\left(\frac{\zeta}{2}|\rho Z^{\rho}_{s}+u^{\rho}_{s}|^{2}+\frac{\zeta(1-\rho^{2})}{2}|Z^{\rho}_{s}|^{2}-a(s,\pi^{\rho}_{s})\right)\mathrm{d}s+Z^{\rho}_{s}\mathrm{d}\bar{B}_{s},\\ &\mathrm{d}\tilde{Y}^{\rho}_{s}=-a(s,\pi^{\rho}_{s})\mathrm{d}s+\tilde{Z}^{\rho}_{s}\mathrm{d}\bar{B}_{s},\\ &u^{\rho}=\sigma\pi^{\rho}=\frac{e^{-\zeta Y^{\rho}}\theta}{(\zeta+1)e^{-\zeta Y^{\rho}}+\gamma},\quad Y_{T}=0,\quad\tilde{Y}_{T}=0.\end{aligned}\right$. (5.1)
Next, we will rigorously verify that the constructed strategy $u^{\rho}$ is an approximate time-consistent equilibrium satisfying Definition 5.1.
Let $R_{t}^{\rho}$ denote the log return at $t$ under strategy $u^{\rho}$. We first show that the Markovian BSDE system (5.1) admits a solution such that $Y^{\rho}$ and $Z^{\rho}$ are uniformly bounded with bounds independent of $\rho\in[-1,1]$ To this end, we need to impose an additional assumption as below.
Assumption 5.2.
We assume that
  1. 1.
    $v$ is differentiable w.r.t. $t,x$ and $v_{x}$ is uniformly bounded.
  2. 2.
    There exists a positive constant $\beta>0$ such that $v,v_{x},m,r,\theta$ are Hölder continuous w. r. t $x,t$ with exponents $\beta,\beta/2$ respectively for any bounded cylinder $Q_{T}$ of the layer $R_{T}$.
By the nonlinear Feynman-Kac formula, if the PDE
$\mathcal{L}f=\zeta\rho f_{x}vu^{\rho}+\frac{\zeta}{2}f_{x}^{2}v^{2}-r-\theta u^{\rho}+\frac{\zeta+1}{2}(u^{\rho})^{2}$
admits a classical solution with the terminal conditionf(T,x)=1f(T,x)=1. Let $Y_{t}:=f(t,X_{t})$ and $Z_{t}:=f_{x}(t,X_{t})v(t,X_{t})$. Then $(Y,Z)$ is the solution to our BSDE. This is a cauchy problem and we introduce the following classical result for the quasi-linear equations with the principal part in divergence form,
$u_{t}-\frac{d}{dx_{i}}a_{i}(t,x,u,u_{x})+a(t,x,u,u_{x})=0,$
with $u(0,x)=\psi_{0}(x)$ in $R_{T}=[0,T]\times\mathbb{R}^{n}$ and
$a_{ij}(x,t,u,p)\equiv\frac{\partial a_{i}(x,t,u,p)}{\partial p_{j}},\quad A(x,t,u,p)\equiv a(x,t,u,p)-\frac{\partial a_{i}}{\partial u}p_{i}-\frac{\partial a_{i}}{\partial x_{i}}.$
Remark 5.3.
The existence of solution to the BSDE system (5.1) can be investigated in xing_class_2018 and FAN20161511, in which a BMO solution can be established. However, we require the boundedness of $Z$ and $\tilde{Z}$ in our verification theorem (Theorem 5.5). Therefore, we employ the Feynman–Kac representation and resort to a PDE-based analysis. Specifically, we utilize Theorem 5.4 as a technical tool to derive the desired boundedness.
Theorem 5.4.
Suppose that the following conditions hold.
  1. $\mathrm{a)}$
    $\psi_{0}(x)\in H^{2+\beta}(\Omega)$ in any $\Omega\subset\mathbb{R}^{n}$ and $\max_{\mathbb{R}^{n}}|\psi_{0}(x)|<\infty.$
  2. $\mathrm{b)}$
    For $t\in(0,T]$ and arbitrary $x,u,p$, we have
    $\mu_{1}\xi^{2}\leq a_{ij}(t,x,u,p)\xi^{2}\leq\mu_{2}\xi^{2},$ (5.2)
    with positive constants $\mu_{1}>0$ and $\mu>0$ and
    $A(x,t,u,0)u\geq-b_{1}u^{2}-b_{2},\quad\text{with constants}\ b_{1},\,b_{2}\geq 0.$
  3. $\mathrm{c)}$
    For any bounded cylinder $Q_{T}$ of the layer $R_{T}$ and $|u|\leq M$, where $M$ is a constant depending only on $b_{1},b_{2}$ and $\psi_{0}$, it holds that _i=1^n ( — a_i — + — aiu — ) (1 + —p—) + _i, j=1^n — aixj — + —a— μ(1 + —p—)^2. Moreover, the functions $a_{i}(x,t,u,p)$ and $a(x,t,u,p)$ are continuous, and $a_{i}(t,x,u,p)$ is differentiable w.r.t. variables $x,u,p$ for each $i$.
  4. $\mathrm{d)}$
    For any bounded cylinder $Q_{T}$ of the layer $R_{T}$ and $|u|\leq M$, $|p|\leq M_{1}$, where $M_{1}$ is a constant depending only on $M,\mu,\mu_{1},\mu_{2}$ and $\psi$, the function $a_{i},a,\frac{\partial a_{i}}{\partial p_{j}},\frac{\partial a_{i}}{\partial u}$, and $\frac{\partial a_{i}}{\partial x_{i}}$ are continuous functions satisfying a Hölder condition in $x,t,u$ and $p$ with exponents $\beta,\beta/2,\beta$ and $\beta$ respectively.
Then there exists at least one solution $u(x,t)$ of the Cauchy problem(5.2) in the strip $R_{T}$ that $|u|\leq M$, $|u_{x}|\leq M_{1}$ and belongs to $H^{2+\beta,1+\beta/2}(\overline{Q_{T}})$ for any bounded cylinder $Q_{T}\subset R_{T}$. It will be an element of $H^{2+\beta,1+\beta/2}(R_{T})$ if it is further assumed that the constants in conditions c)-d) do not depend on $Q_{T}$.
Proof.
The result follows directly from Theorem 6.1 and Theorem 8.1 in ladyzhenskaia1968linear. The dependence of $M$ and $M_{1}$ can be found in Theorem 2.9 and (6.10) in ladyzhenskaia1968linear. ∎
Next, we verify that our equations satisfy these conditions under Assumptions 4.1, 4.2 and 5.2. We reverse the time and still denote f(Tt,x)f(T-t,x) by $f(t,x)$ and rewrite the PDE in the divergence form
$f_{t}-\frac{\mathrm{d}}{\mathrm{d}x}\left(\frac{1}{2}v^{2}(t,x)f_{x}\right)+a(t,x,f,f_{x})=0,$ (5.3)
with $f(0,x)=1$. Here we suppress the dependence of the functions on $(t,x)$ for notational simplicity and also denote that
$a(t,x,f,p)=-mp+\zeta\rho pv\frac{e^{-\zeta f}\theta}{(\zeta+1)e^{-\zeta f}+\gamma}+\frac{\zeta}{2}p^{2}v^{2}-r-\theta\frac{e^{-\zeta f}\theta}{(\zeta+1)e^{-\zeta f}+\gamma}+\left(\frac{e^{-\zeta f}\theta}{(\zeta+1)e^{-\zeta f}+\gamma}\right)^{2}+vv_{x}p,$
$A(t,x,f,p)=-mp+\zeta\rho pv\frac{e^{-\zeta f}\theta}{(\zeta+1)e^{-\zeta f}+\gamma}+\frac{\zeta}{2}p^{2}v^{2}-r-\theta\frac{e^{-\zeta f}\theta}{(\zeta+1)e^{-\zeta f}+\gamma}+\left(\frac{e^{-\zeta f}\theta}{(\zeta+1)e^{-\zeta f}+\gamma}\right)^{2}.$
[…]
and the first two conditions are verified given the differentiability of $v$. For condition c), we have
$\left(\left|a_{i}\right|+\left|\frac{\partial a_{i}}{\partial f}\right|\right)(1+|p|)+\left|\frac{\partial a_{i}}{\partial x_{j}}\right|+|a|=\frac{1}{2}v^{2}|p|(1+|p|)+|vv_{x}p|+|a(t,x,f,p)|\leq C(1+|p|^{2}),$
as $\theta,v,v_{x},m,u^{\rho}$ are bounded independent of $\rho$ and $0\leq\rho\leq 1$. We emphasize that $C$ can be chosen independently of $\rho$. To verify the final condition in the theorem, we employ a minor result concerning Hölder continuous functions. That is, if $h_{1}$ and $h_{2}$ are two bounded Hölder continuous functions with $\alpha_{1}$ and $\alpha_{2}$ being their Hölder coefficients respectively, then $h_{1}h_{2}$ is a Hölder continuous function with Hölder coefficient $\min(\alpha_{1},\alpha_{2})$. Then condition d) can be easily verified and there exists a solution $f\in C^{2+\beta,1+\beta/2}(Q_{T})$ for any bounded cylinder $Q_{T}$ of the layer $R_{T}$. Moreover, as $b_{1},b_{2},\mu,\mu_{1},\mu_{2}$ are independent of $\rho$, the constant $M$ and $M_{1}$ also do not depend on $\rho$. Then $Z^{\rho}=mf_{x}$ is bounded independently of $\rho$.
For the same reason, we seek the solution to
$\mathcal{L}g=-r-\theta\frac{e^{-\zeta f}\theta}{(\zeta+1)e^{-\zeta f}+\gamma}+\frac{1}{2}\left(\frac{e^{-\zeta f}\theta}{(\zeta+1)e^{-\zeta f}+\gamma}\right)^{2},$
with the terminal condition $g(T,x)=0$. Here $f$ is the solution to equation (5.3). By the same argument, $\tilde{Z}^{\rho}$ is also bounded independently of $\rho$.
Next, we turn to show that $\pi^{\rho}$ is an approximate time-consistent equilibrium. Thanks to the concavity of the utility function, it holds that
$J(t,R_{t}^{\rho},\pi^{t,\varepsilon})-J(t,R_{t}^{\rho},\pi^{\rho})\leq e^{R_{t}^{\rho}}\mathbb{E}_{t}\left[U^{\prime}(R_{T}^{\rho})\xi_{T}^{t,\varepsilon}\right]+\frac{\gamma}{2}\left(\textup{Var}(R_{T}^{\rho})-\textup{Var}(R_{T}^{t,\varepsilon})\right)$
Here, we use $\pi^{t,\varepsilon}=\pi^{\rho}+\eta\mathds{1}_{s\in[t,t+\varepsilon)}$ to denote the perturbation strategy and let $\xi^{t,\varepsilon}$ be the resulting difference process for any $\eta\in L^{\infty}_{\mathcal{F}_{t}}(\Omega;\mathbb{R})$ and we shall omit the dependence on $\rho$ and $\eta$ for simplicity.
Theorem 5.5.
Under Assumptions 4.1, 4.2, 5.2, the strategy $\pi^{\rho}$ is admissible for any fixed $\rho\in[0,1]$. Moreover, $\pi^{\rho}$ is an approximate time-consistent equilibrium with the approximate error of the order $O(\rho^{2})$ such that for any $t\in[0,T]$ and any fixed $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega;\mathbb{R})$,
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\left(J(t,R_{t}^{\rho},\pi^{t,\varepsilon})-J(t,R_{t}^{\rho},\pi^{\rho})\right)\leq C\rho^{2}.$
Here $C$ is a positive constant independent of $t$, $\eta$ and $\rho$.
Proof.
As $\pi^{\rho}$ is a bounded strategy, it follows from the same arguments for Theorem 4.4 that $\pi^{\rho}$ is an admissible strategy. We next show that $\pi^{\rho}$ is an approximate Nash equilibrium strategy. Fix $t\in[0,T)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega;\mathbb{R})$ and use $\pi^{t,\varepsilon}$ to denote the perturbed strategy. By Lemma 3.4 and the same argument in Theorem 4.11, it holds that, a.s.,
$J(t,R_{t}^{\rho},\pi^{t,\varepsilon})-J(t,R_{t}^{\rho},\pi^{\rho})$
$\leq$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta Y^{\rho}_{s}}\left((\mu_{s}-r_{s})\eta-\frac{1}{2}\sigma_{s}^{2}\eta^{2}-(\zeta+1)\sigma_{s}\eta u^{\rho}_{s}-\zeta\rho\sigma_{s}\eta Z^{\rho}_{s}\right)\mathrm{d}s\right]-\frac{\gamma}{2}\mathbb{E}\left[\sigma_{s}\eta(2u^{\rho}_{s}+2\rho\tilde{Z}^{\rho}_{s}+\sigma_{s}\eta)\mathrm{d}s\right]$
By the same arguments for Theorem 4.4 and Theorem 4.11, we deduce that
$\lim\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left|\left(e^{-\zeta\left(Y^{\rho}_{s}+R^{\rho}_{s}-R^{\rho}_{t}\right)}-e^{-\zeta Y^{\rho}_{s}}\right)\left((\mu_{s}-r_{s})\eta-\frac{1}{2}\sigma_{s}^{2}\eta^{2}-(\zeta+1)\sigma_{s}^{2}\pi^{\rho}_{s}\eta-\zeta\rho\sigma_{s}\eta Z^{\rho}_{s}\right)\right|\mathrm{d}s\right]=0,\quad a.s..$
Consequently, it holds that
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\left(J(t,R_{t}^{\rho},\pi^{t,\varepsilon})-J(t,R_{t}^{\rho},\pi^{\rho})\right)$
$\leq$ $\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\bigg\{e^{-\zeta Y^{\rho}_{s}}\left((\mu_{s}-r_{s})\eta-\frac{1}{2}\sigma_{s}^{2}\eta^{2}-(\zeta+1)\sigma_{s}\eta u^{\rho}_{s}-\zeta\rho\sigma_{s}\eta Z^{\rho}_{s}\right)\right$.
$\left.-\frac{\gamma}{2}\sigma_{s}^{2}\eta^{2}-\gamma\sigma_{s}\eta(u^{\rho}_{s}+\rho\tilde{Z}^{\rho}_{s})\bigg\}\mathrm{d}s\right]$
$=$ $\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left(-\frac{1}{2}\left(e^{-\zeta Y^{\rho}_{s}}+\gamma\right)(\sigma_{s}\eta)^{2}-\sigma_{s}\eta\left[\zeta e^{-\zeta Y^{\rho}_{s}}\rho Z^{\rho}_{s}+\gamma\rho\tilde{Z}^{\rho}_{s}\right]\right)\mathrm{d}s\right]$
$\leq$ $\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\frac{(\zeta e^{-\zeta Y^{\rho}_{s}}Z^{\rho}_{s}+\gamma\tilde{Z}^{\rho}_{s})^{2}}{2(e^{-\zeta Y^{\rho}_{s}}+\gamma)}\mathrm{d}s\right]\rho^{2}\leq C\rho^{2}.$
Here $C$ is a constant independent of $\rho$ and $t$. Thus $\pi^{\rho}=\sigma^{-1}\frac{e^{-\zeta Y^{\rho}}\theta}{(\zeta+1)e^{-\zeta Y^{\rho}}+\gamma}$ is an approximate time-consistent equilibrium with the desired approximation error $C\rho^{2}$. ∎
6 Numerical Examples
In this section, we present some numerical studies on the equilibrium strategy in the incomplete factor market model by employing some deep learning algorithms developed in e_deep_2017 and Han_2018 to our BSDE systems (4.4), (4.12), and (5.1). In particular, to fulfill our model assumptions, we consider a truncated time-varying Gaussian mean-return model in which the stock price process $S_{t}$ and the market factor process $X_{t}$ evolve as
$\frac{\mathrm{d}S_{t}}{S_{t}}$ $=(r+\sigma\cdot\theta(X_{t}))\mathrm{d}t+\sigma\mathrm{d}B_{t},$
[…]
Here, we choose $\theta(X):=X^{+}\wedge 10000$ to ensure that Assumptions 4.1, 4.2, and 5.2 are satisfied. Similar to dai_dynamic_2021, we fix the parameter values: $r=0.017$, $\sigma=0.15$, $X_{0}=\bar{X}=0.273$, $\lambda=0.27$ and $\nu=0.065$. Moreover, we set $\zeta=1$, $\gamma=0.1$ and $T=2$ for the base model. We discretize the horizon $[0,T]$ into time grid $0=t_{0}<t_{1}<\cdots<t_{N}=T$, where $t_{i}=i\Delta t$ for $i=0,1,\dots,N$ and $\Delta t=T/N$. Henceforth, for notational simplicity, we use the subscript $n$ to denote the processes evaluated at time $t_{n}$. The truncated Ornstein–Uhlenbeck process $X$ is approximated by the Euler scheme:
$X_{n+1}=X_{n}+\lambda(-10000\vee(\bar{X}-X_{n})\wedge 10000)\Delta t+\nu(\bar{B}_{n+1}-\bar{B}_{n})$
for $n=1,\cdots,N-1$, with $X_{0}=0.273$. To numerically illustrate our theoreticla results, we trained three models. For the decoupled BSDE systems (4.4) and (5.1), because the equilibrium strategy $\hat{\pi}$ (for $\rho=0$) and the approximate Nash equilibrium strategy $\pi^{\rho}$ both only depend on the process $Y$, we consider the discretized form of $Y$ in each case, respectively:
$Y_{n+1}=Y_{n}+\left(\frac{\zeta}{2}|\hat{u}_{s}|^{2}+\frac{\zeta}{2}|Z_{n}|^{2}-a(t_{n},\hat{\pi}_{n})\right)\Delta t+Z_{n}(\bar{B}_{n+1}-\bar{B}_{n}),$
and
$Y_{n+1}^{\rho}=Y_{n}^{\rho}+\left(\frac{\zeta}{2}|\rho Z_{n}^{\rho}+u^{\rho}_{n}|^{2}+\frac{\zeta(1-\rho^{2})}{2}|Z_{n}|^{2}-a(t_{n},\pi^{\rho}_{n})\right)\Delta t+Z_{n}(\bar{B}_{n+1}-\bar{B}_{n}).$
We then employ a deep learning approximation for the initial conditions $(Y_{0},Z_{0})$ and the sequence $\{Z_{n}\}_{n=1}^{N-1}$, using $\{X_{n}\}_{n=0}^{N-1}$ as inputs. Particularly, we use N1N-1 fully connected feedforward neural networks with bias terms to represent $\{Z_{n}\}_{n=1}^{N-1}$ and two trainable parameters $(Y_{0},Z_{0})$ to represent the initial values of $Y$ and $Z$ at time $t_{0}$. Motivated by the terminal condition $Y_{T}=0$, we define the mean squared loss function as
$Loss=\mathbb{E}\big[|Y_{N}|^{2}\big].$
For the coupled BSDE system (4.12), we consider the following discretized counterparts:
$\left\{\begin{aligned} &Y_{n+1}=Y_{n}+\left(\frac{\zeta}{2}|\rho Z_{n}+\hat{u}_{n}|^{2}+\frac{\zeta(1-\rho^{2})}{2}|Z_{n}|^{2}-a(t_{n},\hat{\pi}_{n})\right)\Delta t+Z_{n}(\bar{B}_{n+1}-\bar{B}_{n}),\\ &\tilde{Y}_{n+1}=\tilde{Y}_{n}-a(t_{n},\hat{\pi}_{n})\Delta t+\tilde{Z}_{n}(\bar{B}_{n+1}-\bar{B}_{n}).\end{aligned}\right$.
We adopt a similar neural network structure as above, introducing two additional parameters $(\tilde{Y}_{0},\tilde{Z}_{0})$ to represent the initial values of the second BSDE. Moreover, the loss function is modified to $\mathbb{E}\left[Y_{T}^{2}+\tilde{Y}_{T}^{2}\right]$.
Figure 1: The training loss function of an experiment with $\rho=0$.
Note that each neural network has the same structure: one input layer, two hidden layers, and one output layer. Both the input and output layers are 1 dimensional, while each hidden layer has 11 dimensions. We adopt batch normalization immediately after each matrix multiplication and before applying the ReLU activation function. At each epoch, we use the Adam optimizer to update the parameters with mini-batches of $512$ samples. Figure 1 plots the mean of the loss function for the model with $\rho=0$ in the experiment. It is observed that the average loss decreases to approximately $3\times 10^{-5}$ after 5,000 epochs. [^6] Given the small variation range of the learned $Y_{t}$ (see Figure 2), a sufficiently low loss is required for good relative accuracy. We employ a four-stage linear decay learning rate schedule to balance exploration and convergence: (1) exploration: $8\times 10^{-4}\to 5\times 10^{-4}$; (2) convergence: $5\times 10^{-4}\to 2\times 10^{-4}$; (3) fine-tuning: $2\times 10^{-4}\to 5\times 10^{-5}$; (4) terminal reinforcement: $5\times 10^{-5}\to 1\times 10^{-5}$.
Figure 2: A sample trajectory of $Y$, $\hat{\pi}_{t}$ and $X_{t}$ with $\rho=0$
Figure 2 presents a sample trajectory of the learned solution $Y_{t}$, the associated equilibrium strategy $\hat{\pi}_{t}$ and the factor process $X_{t}$. It is observed that the sample path of $Y$ exhibit an overall downward trend over time. In our framework, the process $Y$ satisfies
$Y_{s}$ $=-\frac{1}{\zeta}\log(e^{-\zeta Y_{s}})=-\frac{1}{\zeta}\log\left(\mathbb{E}_{s}\left[e^{-\zeta(R_{T}-R_{s})}\right]\right)$
$=-\frac{1}{\zeta}\log\left(-\mathbb{E}_{s}\left[U(R_{T}-R_{s})\right]\right)=-\frac{1}{\zeta}\log\left(\mathbb{E}_{s}\left[\frac{1}{\zeta}U^{\prime}(R_{T}-R_{s})\right]\right).$
Intuitively, $\mathbb{E}_{s}\left[U(R_{T}-R_{s})\right]$ decreases as time evolves, leading to a decrease in $Y_{s}$. Therefore, $\mathbb{E}_{s}\left[U^{\prime}(R_{T}-R_{s})\right]$ increases as $Y_{s}$ declines. As the investment horizon approaches, the expected marginal utility rises; consequently, the investor becomes more inclined to invest. This also explains why the equilibrium strategy $\hat{\pi}$ in the mean–variance problem in dai_dynamic_2021 remains constant when $\rho=0$, regardless of the volatility of the factor $X_{t}$, since the expected marginal utility is constant in that case.
Figure 3: Plots of the mean of the equilibrium strategy $\hat{\pi}$ of $1000$ trajectories under different $\gamma$ (left panel) and different $\zeta$ (right panel) with $\rho=0$.
Figure 3 plots the mean of the strategy $\hat{\pi}$ using $1000$ trajectories under different values of $\gamma$ and $\zeta$. As illustrated, a higher $\gamma$ corresponds to a smaller investment proportion $\hat{\pi}$, which is intuitively reasonable. This behavior reflects that a larger $\gamma$ indicates the agent places a higher weight on the variance term, leading the investor to favor stability over high returns. Moreover, a higher $\zeta$ also leads to a smaller investment proportion $\hat{\pi}$, which is also intuitively reasonable as $\gamma$ represents the constant risk aversion in the CARA utility function.
Figure 4: Comparison of conditional variance and conditional expected utility: our model with different $\gamma$ vs. model without variance term
Figure 4 compares the conditional variance and expected utility under the equilibrium strategy $\hat{\pi}$ in our model with $\rho=0$ and the corresponding strategy in the same market setting where the agent ignores the variance term. The optimal strategy in this case can be obtained by setting $\gamma=0$:
$\pi=\sigma^{-1}\cdot\frac{\theta(X_{t})}{2}.$
We then simulate 1,000 trajectories of the log-return process under each strategy and compute the sample mean and variance of $R_{T}-R_{t}$ at any time $t=n\Delta t$, $n=0,1,2,\dots,N$. It can be observed that, by incorporating the variance term in the $t$-functional, our equilibrium strategy reduces the conditional variance at the cost of a very small decrease in conditional expected utility.
Figure 5: The mean of $\hat{\pi}_{t}$ and $Y_{t}$ under trading constraint with different $\rho$.
Figure 5 illustrates the equilibrium strategies with trading constraints A=[10000,10000]A=[-10000,10000] and negative correlations ($\rho=-0.31,-0.62,-0.93$) alongside the equilibrium strategy with $\rho=0$. Although the differences in $Y_{t}$ are small, there is a clear gap between the strategies. The constrained cases with negative $\rho$ exhibit higher investment levels. Furthermore, the investment decreases as the terminal time approaches, which is consistent with the results in dai_dynamic_2021.
Time Approximation ($\hat{\pi}^{\rho}$) Constraint ($\hat{\pi}$) Relative Error (‰)
0.0 -0.959920 -0.960158 0.248‰
0.5 -0.968967 -0.969083 0.121‰
1.0 -0.977834 -0.977824 -0.010‰
1.5 -0.990868 -0.990840 -0.028‰
2.0 -1.000000 -1.000000 -0.000‰
Table 1: Comparison of Approximation and Constrained Equilibrium Strategies
Table 1 compares the approximate Nash equilibrium strategy $\hat{\pi}^{\rho}$ with the equilibrium strategy $\hat{\pi}$ under a trading constraint at five discrete time point (t=0.0, 0.5, 1.0, 1.5, 2.0)(t=0.0,\;0.5,\;1.0,\;1.5,\;2.0) with $\rho=-0.31$ and $\gamma=1$ fixed [^7] In this case, since the loss from the variance term is relatively small, we use $\gamma=1$ instead of $0.1$ to make the numerical results more intuitive. . The last column reports the relative errors in per mille (‰). The results show that the discrepancies between the two strategies are negligible, and the relative errors remain extremely small across all time points. Furthermore, as $t\rightarrow T$, the relative error exhibits a slightly decreasing trend. This indicates that the expected payoff generated by the approximate Nash equilibrium strategy closely matches that of the true constrained equilibrium.
Acknowledgements: Zongxia Liang is supported by the National Natural Science Foundation of China under grant no. 12271290. Sheng Wang acknowledges Professor Ka Chun Cheung and the financial supports as a postdoctoral fellow from Department of Statistics and Actuarial Science, School of Computing and Data Science, The University of Hong Kong. Part of this work was completed in 2024, when Sheng Wang was a visiting student (research assistant) under the supervision of Professor Xiang Yu. He acknowledges the financial support under the Hong Kong Polytechnic University research grant under no. P0045654. Xiang Yu is supported by the Hong Kong RGC General Research Fund (GRF) under grant no. 15211524, the Hong Kong Polytechnic University research grant under no. P0045654 and the Research Centre for Quantitative Finance at the Hong Kong Polytechnic University under grant no. P0042708.
Appendix A The Energy Inequality and the John-Nirenberg Inequality
For the reader’s convenience, we collect below the energy inequality and the John-Nirenberg inequality (see kazamaki_continuous_1994 and zhang_backward_2017).
Lemma A.1.
Assume that $Z\in H_{\textup{BMO}}$. For each integer $n\geq 1$, it holds that
$\mathbb{E}\left[\left(\int_{0}^{T}Z^{2}_{s}\mathrm{d}s\right)^{n}\right]\leq n!\|Z\|_{\textup{BMO}}^{2n}.$
Moreover, the conditional version of the energy inequality also holds that, for any stopping time $\tau$,
$\mathbb{E}_{\tau}\left[\left(\int_{\tau}^{T}Z^{2}_{s}\mathrm{d}s\right)^{n}\right]\leq n!\|Z\|_{\textup{BMO}}^{2n},\quad a.s..$
If further $\|Z\|_{\textup{BMO}}<1$, we have the John-Nirenberg inequality.
Lemma A.2.
If $\|Z\|_{\textup{BMO}}<1$, for every stopping time $\tau$
$\mathbb{E}_{\tau}\left[\exp\left(\int_{\tau}^{T}Z_{s}^{2}\mathrm{d}s\right)\right]\leq\frac{1}{1-\|Z\|_{\textup{BMO}}^{2}},\quad a.s..$
Appendix B Sufficient Conditions for the Existence of Solution to BSDEs
For the reader’s convenience, we present a sufficient condition in xing_class_2018 for the existence of a solution to the following BSDE system.
$dY_{t}=-f(t,X_{t},Y_{t},Z_{t})\,dt+Z_{t}\sigma(t,X_{t})\,dW_{t},\qquad Y_{T}=g(X_{T}),$ (B.1)
where $f:[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{N}\times\mathbb{R}^{N\times d}\to\mathbb{R}^{N},\;g:\mathbb{R}^{d}\to\mathbb{R}^{N}$ are Borel functions. We first introduce two structural conditions on the generator $f$ that appear in the theorem.
Definition B.1 (A priori boundedness (AB) condition).
We say that a function
$f:[0,T]\times\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}^{N\times d}\to\mathbb{R}^{N}$
satisfies the condition AB if there exist a deterministic function $l\in L^{1}([0,T])$ and a set of vectors $a_{1},\dots,a_{K}\in\mathbb{R}^{N}$ positively spanning $\mathbb{R}^{N}$, such that
$a_{k}^{\top}f(t,x,y,z)\leq l(t)+\frac{1}{2}\lvert a_{k}^{\top}z\rvert^{2},\qquad\text{for all }(t,x,y,z)\text{ and }k=1,\dots,K.$ (B.2)
We say that $f$ satisfies the weak AB condition (short as $(wAB)$) if there exist Borel functions
$L_{k}:[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{N\times d}\to\mathbb{R}^{d},\quad k=1,\dots,K,$
[…]
$a_{k}^{\top}f(t,x,y,z)\leq l(t)+\frac{1}{2}\lvert a_{k}^{\top}z\rvert^{2}+a_{k}^{\top}z\,L_{k}(t,x,z),\qquad\text{for all }(t,x,y,z)\text{ and }k=1,\dots,K.$ (B.3)
Remark B.2.
A set of nonzero vectors $a_{1},\dots,a_{K}$ in $\mathbb{R}^{N}$ (with $K>N$) is said to positively span $\mathbb{R}^{N}$ if, for each $a\in\mathbb{R}^{N}$, there exist nonnegative constants $\lambda_{1},\dots,\lambda_{K}$ such that
$\lambda_{1}a_{1}+\cdots+\lambda_{K}a_{K}=a.$
Definition B.3 (The Bensoussan–Frehse (BF) condition).
We say that a continuous function
$f:[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{N}\times\mathbb{R}^{N\times d}\to\mathbb{R}^{N}$
satisfies the BF condition if it admits a decomposition of the form
$f(t,x,y,z)=\mathrm{diag}(z\,l(t,x,y,z))+q(t,x,y,z)+s(t,x,y,z)+k(t,x),$ (B.4)
such that the functions
$l:[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{N}\times\mathbb{R}^{N\times d}\to\mathbb{R}^{d\times N}\quad\text{and}\quad q,s,k:[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{N}\times\mathbb{R}^{N\times d}\to\mathbb{R}^{N}$
fulfill the following conditions: there exist $b_{0}\in\mathbb{R}^{d}$ and two sequences $\{C_{n}\}$ and $\{q_{n}\}$ of positive constants with $q_{n}>1+d/2$, and a sequence $\{\kappa_{n}\}$ of functions
$\kappa_{n}:[0,\infty)\to[0,\infty),\qquad\lim_{w\to\infty}\frac{\kappa_{n}(w)}{w^{2}}=0,$
such that for each $n\in\mathbb{N}$ and any $(t,x,y,z)\in[0,T]\times B_{n}(b_{0})\times\mathbb{R}^{N}\times\mathbb{R}^{N\times d}$, we have
$|l(t,x,y,z)|$ $\leq C_{n}(1+|z|),$ (quadratic-linear)
$|q^{i}(t,x,y,z)|$ $\leq C_{n}\left(1+\sum_{j=1}^{i}|z^{j}|^{2}\right),$ $i=1,\dots,N,\qquad\textnormal{(quadratic-triangular)}$
$|s(t,x,y,z)|$ $\leq\kappa_{n}(|z|),$ (subquadratic)
$k$ $\in L^{q_{n}}([0,T]\times B_{n}),$ $\textnormal{(z-independent)}.$
In this case, we write $f\in\mathrm{BF}(\{C_{n}\},\{\kappa_{n}\},\{q_{n}\})$.
Theorem B.4 (Existence under BF + AB conditions).
Suppose that $f$ satisfies BF and AB conditions, and that $g\in C^{\{\alpha_{n}\}}_{\mathrm{loc},b_{0}}$ for some $b_{0}$ and it satisfies $\lim_{|x|\to\infty}\frac{|g(x)|}{|x|^{2}}=0$. Then the system (B.1) admits a locally Hölderian solution $(v,w)$, that is, $v\in C^{\{\alpha_{n}^{\prime}\}}_{\mathrm{loc},b_{0}}$ for some sequence $\{\alpha_{n}^{\prime}\}$ in $(0,1]$. When $g$ is bounded, the AB condition can be replaced by wAB condition and $(v,w)$ is a bounded bmo-solution.
Appendix C The Challenge of our BSDE System when $\rho\neq 0$
For the general case $\rho\neq 0$, we encounter a system of coupled two-dimensional quadratic BSDEs. Existing results for the existence of solution to the multidimensional BSDEs bifurcate into two paradigms: fixed-point arguments via contraction mappings (yielding the unique solution) (see FAN2023105, Luo_2017 and HU20161066) and construction of uniformly convergent approximants through coefficient regularization (see xing_class_2018, Jackson_2022 and JACKSON2023). In this section, we explain that these two methods are not applicable in our setting.
[…]
$\frac{\zeta}{2}z^{2}-\rho\zeta z\cdot\frac{\zeta e^{-\zeta y}\rho z+\gamma\rho\tilde{z}}{(\zeta+1)e^{-\zeta y}+\gamma}+\frac{\zeta+1}{2}\left(\frac{\zeta e^{-\zeta y}\rho z+\gamma\rho\tilde{z}}{(\zeta+1)e^{-\zeta y}+\gamma}\right)^{2}$
[…]
$\frac{1}{2}\left(\frac{\zeta e^{-\zeta y}\rho z+\gamma\rho\tilde{z}}{(\zeta+1)e^{-\zeta y}+\gamma}\right)^{2}.$
The first approach typically requires a continuity condition on $y$ to guarantee a contraction mapping that
[…]
which is not satisfied in our case as $y$ is coupled with the quadratic term of $z,\tilde{z}$.
The second approach relies on the AB condition or the wAB condition (see Appendix B) for the convergence of a subsequence of approximated solutions to the approximated BSDE system. Specifically, there exist a deterministic function $l\in L^{1}[0,T]$ and $a_{1}\cdots a_{k}$ that positively span $\mathbb{R}^{N}$ such that
$a_{k}^{\top}f(t,x,y,z)\leq l(t)+\frac{1}{2}|a_{k}^{\top}z|^{2}+a_{k}^{\top}z\,L_{k}(t,x,z).$
[…]
Here, $L_{i}(t,x,z,\tilde{z})\leq C(1+|z|+|\tilde{z}|),\;i=1,2$ for a constant $C$.
[…]
$\frac{\zeta e^{-\zeta y}\rho z+\gamma\rho\tilde{z}}{(\zeta+1)e^{-\zeta y}+\gamma}\longrightarrow\rho\tilde{z},\quad\frac{\zeta e^{-\zeta y}\rho z+\gamma\rho\tilde{z}}{(\zeta+1)e^{-\zeta y}+\gamma}\longrightarrow\frac{\zeta\rho z}{\zeta+1}.$
Substituting it into wAB condition, we obtain
$\frac{\zeta}{2}b^{2}+-\left(\rho^{2}\zeta+\frac{\rho^{2}}{2}\right)b+\frac{\zeta+1}{2}\rho^{2}\leq 0,\quad\left(\frac{\zeta}{2}-\frac{\zeta^{2}\rho^{2}}{2(\zeta+1)}\right)b^{2}-\frac{\rho^{2}\zeta^{2}}{2(\zeta+1)^{2}}b^{3}\leq 0.$ (C.1)
and
$\frac{1}{2}\rho^{2}c^{2}-c\left(\frac{\zeta}{2}-\rho^{2}\zeta c+\frac{\zeta+1}{2}\rho^{2}c^{2}\right)\leq 0,\quad\frac{\zeta^{2}\rho^{2}}{2(\zeta+1)^{2}}-c\left(\frac{\zeta}{2}-\frac{\zeta^{2}\rho^{2}}{2(\zeta+1)}\right)\leq 0.$ (C.2)
It then follows from (C.1) that
$\rho^{4}(\zeta+\frac{1}{2})^{2}-\zeta(\zeta+1)\geq 0,\quad\frac{\zeta+1}{\zeta}\leq\frac{(\zeta+1)^{2}}{\rho\zeta}-(\zeta+1)\leq b\leq\frac{\rho^{2}\zeta+\frac{\rho^{2}}{2}+\sqrt{\rho^{4}(\zeta+\frac{1}{2})^{2}-\zeta(\zeta+1)}}{\zeta}\leq\frac{\zeta+1}{\zeta},$
which result in $\rho=1$ and $b=\frac{\zeta+1}{\zeta}>0$. Substituting $\rho=1$ into (C.2), we get that $c>0$ by (C.1). Consequently, we have
$\frac{\zeta+1}{2}c^{2}-\left(\zeta+\frac{1}{2}\right)c+\frac{\zeta}{2}\geq 0,\quad c\geq\frac{\zeta}{\zeta+1},$
which then imply $c\geq 1$.
As we have shown that, for any $a_{i}=(m,n)\in\{a_{1},\cdots,a_{k}\}$ chosen in wAB condition, $m$ and $n$ must satisfy one of the three conditions
  1. 1.
    $m\geq 0$, $n\geq 0$.
  2. 2.
    $m<0$, $n=-\frac{\zeta+1}{\zeta}m$.
  3. 3.
    $n<0$, $m\geq n$.
Each of the three cases implies that $(m,n)\cdot(-1,-1)\leq 0$, as a consequence, $\{a_{1},\cdots,a_{k}\}$ cannot positively span $\mathbb{R}^{2}$. Then neither AB condition nor wAB condition holds in our problem.